### v1

# Using Biot-Savart to shrink Eulerian domains while maintaining or improving external flow accuracy

Abstract
We introduce a novel boundary condition formulated using a Biot-Savart vorticity integral inside an Eulerian fluid domain that maintains high-accuracy results even when the domain boundary is within a body-length of immersed solid boundaries. The Biot-Savart condition is incorporated into a general solver with Eulerian velocity-pressure fields by including the pressure-induced body vorticity during the incompressible-flow projection-step. We use a multilevel approach to reduce the computational cost of the evaluation of the Biot-Savart integral from $O(N)$ to $O(\log N)$ and formally show that the error associated with this approach is bounded by the total vorticity contained in the lower levels and that it scales with the inverse power of the problem’s dimension. We use two-dimensional and three-dimensional analytic vortex problems to demonstrate this scaling is correct and select appropriate multilevel coarsening parameters accordingly. We show that the new method exactly captures the analytical added-mass force of a square plate accelerated from rest even when the entire domain only extends 1/2 diameter from the plate. The new method also predicts essentially identical time-varying drag forces for a two-dimensional circle regardless of domain size, while classical boundary conditions require a domain nearly 10 times larger to converge on the new method’s result. Finally, we study the sensitivity of deflected wakes to the new domain boundary conditions and show that only in minimal domains, 5 times smaller than the reference domain, the omission of the far-field wake influences the flow field and the forces on the body. Doubling the domain size recovers the correct results.
keywords:
Biot-Savart , Boundary Condition , Projection Method , Unbounded Domains
PACS:
$0$ , $1$
MSC:
$0$ , $1$
journal: Journal of Computational Physics
[…]
[inst1]organization=Ship Hydromechanics,addressline=Faculty of Mechanical Engineering (mE),
Delft University of Technology, city=Delft, country=Netherlands
{graphicalabstract}[Uncaptioned image]
{highlights}
External flow boundary condition via the Biot-Savart integral enables minimal computational domains
Fast ($\log N$) multilevel evaluation of the Biot-Savart integral with controllable bounded error.
Reduction in degrees of freedom by $O(10^{3})$ compared to reflection conditions for flow with bounded vorticity.
1 Introduction
Computational fluid dynamics provides a numerical solution to a continuum problem in an (often) unbounded domain. Common practice is to approximate the continuum problem with a discrete set of equations and truncate the unbounded domain to a finite computational domain. While the former does not usually introduce significant errors, the latter can significantly impact the numerical solution when the finite domain is not sufficiently large. In the context of external flows, sufficiently large means that the decaying potential flow induced by the body and the advection/diffusion of vorticity through the domain’s boundary have no influence the forces experienced by the body [1].
Numerical simulations of external incompressible flows enforce far-field boundary conditions on the computational domain’s exterior. These consist of a known velocity normal to the boundary and a known normal pressure gradient111for a description of the boundary conditions for the incompressible Navier-Stokes equations, we refer the reader to [2]. These can lead to significant errors. In particular, the force acting on a body can suffer from severe blockage effects when the computational domain is not sufficiently large. While domain sensitivity study can determine the extent of the blockage, there is no a priori rule for estimating their magnitude, making this an expansive iterative process.
Grid stretching away from the immersed body can efficiently maximize the domain size while maintaining a high-resolution mesh around the body, see [3, 4]. This can limit the blockage effects on the body. However, this method requires correct stretching ratios to avoid excessive numerical dissipation or reflection inside the domain. The method is also inefficient in cases where the body undergoes substantial motion within the computational domain, for example, when tracking the motion of a kite. Grid stretching can be replaced by a buffer layer that acts as a numerical damper which ensures that the solution decays rapidly before it reaches the boundaries [5]. Following a similar approach, a coarse grid can provide the boundary conditions for a relatively small, well-resolved domain; this is the basis of the multi-domain methods [1]. The method interpolates the coarse mesh solution to the walls of the finer domain, providing artificial far-field boundary conditions.
For flows with one [6, 7] or two unbounded directions [8], coordinate transforms or analytical potential flow solution can be applied in the unbounded directions and correctly capture solution. Potential flow boundary conditions are prescribed following the classical Hodge decomposition of the velocity field into an irrotational and an incompressible part
$\vec{u}=\vec{u}_{\omega}+\nabla\phi+\vec{U}_{\infty}=-\nabla^{-2}\left(\nabla\times\vec{\omega}\right)+\nabla\phi+\vec{U}_{\infty}$ (1)
where $\vec{u}_{\omega}$, $\nabla\phi$, $\vec{U}_{\infty}$ are the incompressible component, irrotational component, and the free-stream velocity, respectively.
Fundamental solution to discrete operators (Lattice Green’s functions or LGFs) and their asymptotic expansions can be used to approximate discrete differential equations on unbounded domains [9]. Points close to the source are evaluated with the discrete kernel (accurate) while points far away are evaluated with the asymptotic expansion (fast). These methods rely on the unbounded properties of the LGFs and can simulate unbounded viscous incompressible flow [10] and flows with immersed-boundaries [11]. However, they are based on the solution of inhomogeneous, linear, constant coefficient elliptic equations and are thus incompatible with some immersed-boundary methods [3, 4].
Other methods leverage the strength of an Eulerian-based solver for the near body and Lagrangian solvers for the far wake by coupling a near wall Eulerian solver to a Vortex particle method in the wake [12]. They enable tracking the coarser wake for an extended time while enabling accurate modeling of the developing boundary layer on the object’s surface. However, several (user-defined) parameters are required to correctly interpolate the Vortex particle method solution onto the Eulerian mesh. Extension to three-dimensional flow is possible but will significantly increase the computational cost.
All these methods allow for a reduction of the error introduced by truncating an infinite computational domain. However, their generality is often limited to specific discretization of the governing equations or introduces complex coupling procedures.
In this manuscript, we propose a novel method for imposing external potential-flow boundary conditions on the velocity and pressure fields on the exterior of a Cartesian computational domain. The method retains the primitive variables formulation ($\vec{u}$ and $p$) of the covering equations and is compatible with immersed boundaries. Additionally, it does not introduce additional Lagrangian or Eulerian mesh. In the next section, we describe the Biot-Savart boundary conditions and how they are injected into a projection algorithm to solve the Navier-Stokes equations. Next, we demonstrate the accuracy of the velocity reconstruction and the balance between speedup and accuracy in the reconstructed field. We then demonstrate, with confined vorticity and vortex-street examples, that the new boundary conditions significantly outperform standard reflective boundary both in terms of computational time and solution accuracy. We present unexpected result that the new method continues to perform well when the fundamental assumption of potential flow external to the domain is violated, and use the unstable deflected wakes generated by fast flapping foils to study the limits of the approach. We close this manuscript with a discussion of the method’s applicability to various flows in science and engineering.
2 Method
$\mathcal{B}$
$\Omega$
$\partial\mathcal{B}$
$\hat{n}$
$\partial\Omega$
Figure 1: Schematic of the fluid domain $\Omega$ with an immersed body $\mathcal{B}$ and their common (wet) interface $\partial\mathcal{B}$. The exterior of the computational domain $\Omega$ is denoted as $\partial\Omega$ with $\hat{n}$ the unit normal vector.
We start by describing the governing equations and their numerical solution. The flow is governed by the Navier-Stokes equations for an incompressible flow. We discretize these equations on a staggered finite-volume mesh. The resulting system of equations is solved within a computational domain $\Omega(N)$, where $N$ is the number of degrees of freedom of the problem. The outer domain boundary is denoted at $\partial\Omega(N^{\cal S})$, where $\cal S$ is a factor linked to the dimensions of the problem. In 2D we have $S=1/2$, and in 3D $S=2/3$. The primary source of vorticity in the domain is an immersed body $\mathcal{B}$ with outer (wet) surface $\partial\mathcal{B}$ and unit normal vector $\hat{n}$, see Figure 1.
We use the standard projection scheme [13] to solve the coupled velocity-pressure system inside a predictor-correct method to achieve second-order temporal convergence of the flow variables [4]. To simplify the discussion, we present the method for the first step of the predictor-corrector; generalization to the second-order Heun’s corrector method or higher-order time integration method is straightforward. The projection method starts with an explicit estimate of the intermediate velocity field $u^{*}$
$\vec{u}^{*}=\vec{u}\,^{t}+\int_{t}^{t+\Delta t}\frac{1}{Re}\nabla^{2}\vec{u}-\left(\vec{u}\cdot\nabla\right)\vec{u}\text{ d}t$ $\quad\forall\ \vec{x}\in\Omega(N),$ (2)
where $Re$ is the Reynolds number of the flow. Depending on the choice of integration method for the right-hand-side, we obtain explicit or semi-implicit methods. Here we will focus on fully explicit methods. On the immersed body $\cal B$, the intermediate velocity is subject to the condition
$\vec{u}^{*}=\vec{U}_{b}\quad\forall x\in\partial\cal{B}.$ (3)
where $\vec{U}_{b}$ is the body velocity. Body-fitted methods enforce this condition directly, while immersed boundary methods enforce it indirectly; either by reformulating the governing equations near the boundary such as [14, 3] or by modifying the discretization near $\cal{B}$ [15].
Additionally, Eq. (2) must be supplemented with far-field boundary conditions on the domain’s outer boundaries. As discussed in the introduction, the simplest of these is a reflection condition
$\frac{\partial\vec{u}^{*}_{s}}{\partial\hat{n}}=0,\ \vec{u}^{*}_{n}=U_{n}$ $\quad\forall\ \vec{x}\in\partial\Omega(N^{\cal S}),$ (4)
where we have decomposed the velocity field on the boundary in a boundary normal $\vec{u}^{*}_{n}$ and tangential part $\vec{u}^{*}_{s}$. We refer the reader to [2] for a discussion of pressure boundary conditions on the domain’s exterior.
Projection of the divergent velocity field onto the solenoidal space is achieved by computing a pressure field $p$ that removes the divergent part of the intermediate velocity field
$\nabla\cdot\frac{\Delta t}{\rho}\nabla p=\nabla\cdot\vec{u}^{*}$ $\quad\forall\ \vec{x}\in\Omega(N),$ (5)
$\vec{u}^{t+\Delta t}=\vec{u}^{*}-\frac{\Delta t}{\rho}\nabla p$ $\quad\forall\ \vec{x}\in\Omega(N),$ (6)
also subject to the standard boundary condition
$\frac{\partial\vec{u}^{t+\Delta t}_{s}}{\partial\hat{n}}=0,\ \vec{u}^{t+\Delta t}_{n}=U_{n}$ $\quad\forall\ \vec{x}\in\partial\Omega(N^{\cal S}).$ (7)
As mentioned previously, the boundary $\partial\Omega$ must be placed sufficiently far away from the body not to influence the fast decaying (potential flow) part of the velocity field and ensure enough space to properly advect the vorticity downstream.
2.1 Biot-Savart boundary conditions
We propose to replace the typical “local” domain boundary conditions such as Eq. (4) &  (7) by a Biot-Savart integral over the vorticity inside the domain. A Green’s function solution to the vorticity equation Eq. (1) gives the velocity at a point $\vec{x}$ as
${\vec{u}}({\vec{x}})=f(\vec{x};\vec{\omega},\Omega)=\int_{\Omega}K_{n}({\vec{x}}-\vec{y})\times\vec{\omega}({\vec{y}})\text{ d}\vec{y}\quad\quad\forall\vec{x}\,\in\partial\Omega(N^{\cal S})$ (8)
where $K_{n}$ is the $n$-dimensional Biot-Savart kernel. This kernel takes the following form [16]
$2D:\quad K(\vec{r})\equiv-\frac{\vec{r}}{2\pi|\vec{r}|^{2}},\qquad\qquad 3D:\quad K(\vec{r})\equiv-\frac{\vec{r}}{4\pi|\vec{r}|^{3}}.$ (9)
This formulation implicitly assumes the flow external to the domain is irrotational and that the flow is incompressible everywhere. Such a formulation is exact for the flow around accelerated bodies when the time $\frac{tU}{L}\lesssim t^{\text{crit.}}$, with $t^{\text{crit.}}$ the critical time for vortex detachment [17]. However, we will show that the results of using equation (8) in an Eulerian solver are still highly accurate for flows with wakes that extend to the domain boundary.
Note that on a staggered grid, we only need to apply the integral equation (8) to compute the velocity normal to the domain face. For ghost cells external to the computational domain, we use the local derivative conditions $\nabla\cdot\vec{u}=\nabla\times\vec{u}=0$ to compute the velocity. These conditions are consistent with the assumption that the flow is potential outside the domain and fast to apply. Specifically, we apply the zero curl condition to all tangential velocity components in the ghost cells and then the zero divergence condition to the normal components in the ghost cells, marching away from the domain boundary.
2.2 Fast evaluation of the Biot-Savart integral
$\mathcal{D}^{(4)}$
$\mathcal{D}^{(3)}$
$\mathcal{D}^{(1)}$
$\mathcal{D}^{(2)}$
$\partial\Omega$
$\Omega$
$\mathcal{B}$
$\vec{y}\,^{(2)}$
$\vec{y}\,^{(3)}$
$\vec{x}$
${S}\,^{(2)}$
$h^{(4)}$
$h^{(3)}$
${S}\,^{(2)}$
Figure 2: Schematic of the multilevel approach to evaluate the Biot-Savart integral for a boundary point $\vec{x}$ over the nested domains $\mathcal{D}^{(1)}\cup\mathcal{D}^{(2)}\cup\mathcal{D}^{(3)}\cup\mathcal{D}^{(4)}$. The second domain half-width is shown, ${S}^{(2)}$. The computational domain $\Omega$ and its boundary $\partial\Omega$ and the immersed body $\mathcal{B}$ are shown. We also show a schematic of the coarsening of the grid between the domains. Domain and cell size are only for representation proposes.
The aim of this novel boundary conditions is to reduce the computational cost compared to a large grid by providing an accrate estimate of the far field conditions. A naive evaluation of Eq. 9  would require $O(N)$ operations for each boundary point, meaning $O(N^{1+\cal S})$ cost overall, overwhelming any savings from reducing the size of the domain. In vortex-particle methods, the integral is usually evaluated with Fast Multipole Method that allows reducing this cost to $O(\log N)$ operations. Here, we use a multilevel algorithm to achieve the same speed-up in our Eulerian setting.
We construct the multilevel Biot-Savart operator for each point $\vec{x}$ on the domain boundary $\partial\Omega$ by representing the complete domain $\Omega$ with a combination of gradually coarser meshes
$\Omega=\mathcal{D}^{(1)}(\vec{x})\cup\mathcal{D}^{(2)}(\vec{x})\cup\cdots\cup\mathcal{D}^{(l-1)}(\vec{x})\cup\mathcal{D}^{(l)}(\vec{x}),$ (10)
where level $1$ is the finest domain and $l$ is the coarsest, and each domain is non-overlapping and their union fully covers $\Omega$, Fig 2. In this work, we define nested rectangular boxes of half-width $S^{(i)}$ centered on $\vec{x}$ and scale this width with the cell size which doubles at each level, i.e. $h^{(i)}=h^{(1)}2^{i-1}$. Therefore the domains are determined by $\vec{x}$ and a non-dimensional half-width
$\tilde{S}=\frac{{S}^{(l)}}{2^{l-1}},$ (11)
which is equal on every level. This domain decomposition is chosen so that the contributions to Eq. (8) are more accurately calculated when close to $\vec{x}$, and the smaller contributions further from $\vec{x}$ are quickly estimated on the coarser levels.
With this set of $l$ non-overlapping domains $\mathcal{D}^{(i)}$, the integral in Eq. 8 can be replaced with a sum over all the contribution from the domains
$f(\vec{x};\vec{\omega},\Omega)\approx\sum_{i=1}^{l}f(\vec{x};\vec{\omega}^{(i)},\mathcal{D}^{(i)}).$ (12)
Once the multilevel field $\omega^{(i)}$ is computed, the total number of operations to evaluate equation 12 for a simulation of dimension $d=2$ or $3$ is therefore $O(l\tilde{S}^{d})=O(\log N)$ since $\tilde{S}$ is constant and $l=O(\frac{1}{d}\log N)$.
In practise, we use the circulation over a cell face $\Gamma=\omega h^{2}$ instead of the vorticity in the Biot-Savart formulation. The circulation on a coarser level is calculated by uniformly pooling the circulation on the level above
$\Gamma^{(i+1)}=\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}$ (13)
where $\Gamma^{(1)}$ is the original finest level field and the pooling $\mathcal{P}$ for a cell at level $i$ simply sums over the corresponding subcells in level $i-1$, see [18]. Thus, our multi-level field $\Gamma^{(i)}$ conserves the total circulation to machine precision, as recommended by [1] and others.
Finally, we derive the error induced by this multilevel approach, starting with the 2D-case and extending the result to 3D. The velocity at location $\vec{x}$ induced by the circulation from a single cell on level $i+1$ is given by (see Fig. 2)
$\vec{u}^{(i+1)}(\vec{x})=\frac{\vec{r}}{2\pi|\vec{r}|^{2}}\times\vec{e}_{z}\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}$ (14)
where $\vec{r}=\vec{y}-\vec{x}$, $y$ is the center of the cell, and $\vec{e}_{z}$ is the out-of-plane unit vector. The velocity induced by the four corresponding cells on level $i$ is
$\vec{u}^{(i)}(\vec{x})=\mathcal{P}^{(i)\to(i+1)}\left(\frac{\vec{r}_{j}}{2\pi|\vec{r}_{j}|^{2}}\times\vec{e}_{z}\Gamma^{(i)}_{j}\right),$ (15)
where $\vec{r}_{j}=\vec{y}_{j}-\vec{x}$ and $y_{j}$ are the four cell centers. We can further decompose this expression by noting that $\vec{r}_{j}=\vec{y}-\vec{x}-\vec{\delta r}_{j}=\vec{r}-\vec{\delta r}_{j}$. For sufficiently large $\vec{r}$ we can assume that $|\vec{r}|\approx|\vec{r}-\vec{\delta r}_{j}|$ which allows us to separate the kernel’s influence into two parts
$\vec{u}^{(i)}(\vec{x})=\mathcal{P}^{(i)\to(i+1)}\left(\frac{\vec{r}}{2\pi|\vec{r}|^{2}}\times\vec{e}_{z}\Gamma^{(i)}_{j}-\frac{\vec{\delta r}_{j}}{2\pi|\vec{r}|^{2}}\times\vec{e}_{z}\Gamma^{(i)}_{j}\right).$ (16)
After rearrangement, the first term equals Eq. (14). The second term is the error due to the multilevel approximation. Its magnitude is
$\varepsilon(\vec{x})^{(i)}=\mathcal{P}^{(i)\to(i+1)}\frac{|\vec{\delta r}_{j}|}{2\pi|\vec{r}|^{2}}\Gamma^{(i)}_{j},$ (17)
as $\vec{e}_{z}$ is a unit vector. On uniform Cartesian meshes, $|\vec{\delta r}_{j}^{(i)}|\sim 2^{l}$, while $|\vec{r}|^{2}\sim(2^{l}\tilde{S})^{2}$. An upper bound for the total error in the induced velocity at a point $\vec{x}$ due to the $i$th-level evaluation of the Biot-Savart integral is therefore
$\varepsilon(\vec{x})\approx\frac{1}{2\pi\tilde{S}^{2}}\sum_{i=2}^{l}\frac{1}{2^{i}}\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}.$ (18)
which decays as $\tilde{S}^{-2}$. The content of each sum is a decreasing fraction of the total vorticity contained in the domain $\mathcal{D}^{(i)}$. As expected, taking $\tilde{S}>\text{Domain}$ yields $\varepsilon(\vec{x})=0$, and as $\Gamma^{(i)}\to 0,\varepsilon(\vec{x})\to 0$, finally as $\tilde{S}\to 0,\varepsilon(\vec{x})\to\infty$. Similarly, in 3D the upper bound for the error is given by
$\varepsilon(\vec{x})\approx\frac{1}{4\pi\tilde{S}^{3}}\sum_{i=2}^{l}\frac{1}{2^{2i}}\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}.$ (19)
which decays as $\tilde{S}^{-3}$. Similarly, the content of the sum is a (even smaller) decreasing fraction of total vorticity contained in each level, improving the method’s accuracy in 3D compared to 2D.
2.3 Biot-Savart & Projection Algorithm
Substitution of the Biot-Savart boundary condition (Eq. 8) into the projection scheme results in a slightly modified intermediate velocity field
Biot-Savart intermediate step
$\vec{u}^{*}=\vec{u}\,^{t}+\int_{t}^{t+\Delta t}\frac{1}{Re}\nabla^{2}\vec{u}-\left(\vec{u}\cdot\nabla\right)\vec{u}\text{ d}t$ $\quad\forall\ \vec{x}\in\Omega(N),$ (20)
$\vec{u}^{*}=f(\vec{x}_{b},\nabla\times\vec{u}^{*},\Omega)$ $\quad\forall\ \vec{x}_{b}\in\partial\Omega(N^{\cal S})$ (21)
where the exterior boundary condition have been computed from the intermediate velocity $\vec{u}^{*}$ in the domain. The projection step, however, results in a coupling between the body and the domain’s boundary through the pressure field
$\nabla\cdot\frac{\Delta t}{\rho}\nabla p=\nabla\cdot\vec{u}^{*}$ (22)
$\vec{u}^{t+\Delta t}=\vec{u}^{*}-\frac{\Delta t}{\rho}\nabla p$ (23)
$\vec{u}^{t+\Delta t}=f(\vec{x},\nabla\times\vec{u}^{t+\Delta t},\Omega)=\vec{u}^{*}-f(\vec{x}_{b},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega)$ (24)
Where we have omitted the domain of definition of each equation for clarity. Away from the body, the source term in the Biot-Savart integral in Eq. 24  vanishes. However, due to the pressure boundary condition on the immersed body ($\partial p/\partial\hat{n}=\hat{n}\cdot(\nabla^{2}\vec{u}/Re-\text{D}\vec{U}_{b}/\text{D}t)$), a thin layer of vorticity is generated by the pressure gradient term that changes the induced velocity field on the outer domain’s boundary. The projection steps and the Biot-Savart boundary conditions now form a fixed-point system
Biot-Savart projection step
$\nabla\cdot\frac{\Delta t}{\rho}\nabla p=\nabla\cdot\vec{u}^{*}-\nabla\cdot f(\vec{x}_{b},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega)$ (25)
$\vec{u}^{t+\Delta t}=\vec{u}^{*}-\frac{\Delta t}{\rho}\nabla p$ (26)
$\vec{u}^{t+\Delta t}=f(\vec{x},\nabla\times\vec{u}^{t+\Delta t},\Omega)=\vec{u}^{*}-f(\vec{x}_{b},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega)$ (27)
This coupling only occurs on the domain’s boundary, and the resulting system is mildly non-linear. This weak non-linearity can be solved efficiently with a deferred correction approach, which explicitly updates the source term of the Poisson problem during each pressure solver iteration. This coupling is only apparent in immersed boundary methods if the pressure boundary conditions are correctly imposed on the immersed body during the projection step; see for example [19, 4].
Lastly, we note that the small multilevel error discussed in the previous section means that global mass conservation is no longer guaranteed on the domain boundary. As such, we apply a uniform correction
$\vec{u}=\vec{u}-\frac{\oint_{\partial\Omega}\hat{n}\cdot\vec{u}\text{ d}S}{\oint_{\partial\Omega}\text{ d}S}.$ (28)
ensuring a globally divergence-free field is passed to the pressure projection method.
3 Multilevel reconstruction validation
We start by validating the multilevel reconstruction algorithm presented above. We selected to cases such respect our assumption of vorticity being confined far from the domain’s boundary. We demonstrate the 2D and 3D reconstruction error of the multilevel and the speedup generated by selecting the correct kernel radius from the above analysis.
Our Biot-Savart boundary conditions are implemented inside WaterLily.jl [^2] https://github.com/weymouth/WaterLily.jl, a fast, immersed-boundary flow solver based on the Boundary data immersion method [3] that can execute both on CPU and GPU [20]. To leverage the GPU capability of the flow solver and fully benefit from our novel boundary conditions, we perform all computations presented herein on an NVIDIA-A6000 GPU (48GB memory). All the codes used for this study and the instructions to reproduce these results are freely available at https://github.com/weymouth/BiotSavartBCs.jl.
3.1 2D validation: Lamb vortex
The velocity field generated by the Lamb-Chaplygin dipole is obtained from the scalar stream function $\vec{u}=-\vec{e}_{z}\times\nabla\psi$ given by
$\psi(r)=\begin{cases}\frac{-2UJ_{1}(\beta r)}{\beta J_{0}(\beta R)},\quad\text{for}\quad r<R,\\ U\left(\frac{R^{2}}{r}-r\right),\quad\text{for}\quad r\geq R,\end{cases}$ (29)
where $J_{0}$ and $J_{1}$ are the zeroth and first Bessel functions of the first kind, respectively. $\beta$ is a parameter with the value $\beta=1.2197\pi/R$. The core of the vortex is defined by the isoline $\psi(r=R)=0$, that delimits the inside and outside of the vortex.
(a) $\log(\text{size})=1$
(b) $\log(\text{size})=3$
(c) $\log(\text{size})=5$
(d) $\log(\text{size})=8$
Figure 3: $L_{2}$-norm of the error in reconstructed velocity field outside the core of the 2D Lamb-Chaplygin vortex for varying (multi-level) kernel size. Error bars are shown in log-scale.
We reconstruct the velocity field outside the vortex core using the internal vorticity and our multilevel Biot-Savart method and compare that to the analytic result. We reconstruct the velocity on a domain of size $[2^{8},2^{8}]$. The dependency of the absolute error in the reconstructed velocity field on the radius of the multilevel refinement criterion are shown in Fig. 3. Setting the refine size $\tilde{S}$ to the domain size $\cal D$ eliminates the multi-level-error, but the result is not exact since the analytic vorticity distribution is sampled on a finite grid of size $\Delta x=0.05R$. This error depends primarily on the distance from the vortex, Fig. 3(d). Progressively reducing the refinement size $\tilde{S}$ introduces the multilevel error quantified in the previous section, which is only weakly dependent on the distance to the vortex.
This is further quantified in Fig. 4(a), which shows the maximum error in the reconstructed velocity field at a distance $d$ from the vortex from the Lamb-Chaplygin vortex. The size $\tilde{S}$ is shown to dominate the reconstruction error, with fairly uniform total error levels once $\tilde{S}\leq 2^{3}=8$. Fig. 4(a) reports the measured computational speed-up achieved by the multilevel method relative to setting $\tilde{S}=\cal D$, demonstrating up to 300$\times$ acceleration.
Figure 4: Maximum error in the reconstructed velocity field generated by a 2D Lamb-Chaplygin vortex dipole for various kernel sizes and distances from the vortex core (left). Corresponding speed-up obtained by reducing the kernel size for the multilevel algorithm (right). The mesh has dimensions $[2^{8},2^{8}]$.
3.2 3D validation: Hill vortex
This validation is repeated in 3D using the classical Hill vortex, whose radial and tangential velocity components are given by
$v_{r}$ $=\frac{1}{r^{2}\sin\theta}\frac{\partial\psi}{\partial\theta},\qquad v_{\theta}=-\frac{1}{r\sin\theta}\frac{\partial\psi}{\partial r}.$ (30)
and the stream function
$\psi(r,\theta)=\begin{cases}-\frac{3U}{4}\left(1-\frac{r^{2}}{R^{2}}\right)r^{2}sin^{2}\theta\quad\text{for}\,r\leq R,\\ \,\,\frac{U}{2}\left(1-\frac{r^{3}}{R^{3}}\right)r^{2}sin^{2}\theta\quad\text{for}\,r>R.\end{cases}$ (31)
with the corresponding errors and speed-up shown in Figure 5. Note that the analytic vorticity of the hill vortex is discontinuous, increasing the truncation error near the vortex compared to the smooth Lamp-vortex, a problem which would not be observed with any real three-dimensional viscous flow.
As argued in Sec. 2.3 the multilevel error decreases extremely quickly as $\tilde{S}$ is increased for three-dimensional flows. This is due to the scaling of the Biot-Savart kernel itself. The 3D case demonstrates a stronger sensitivity to the distance to vortex, as expected from the error estimate. Similarly, the speed-up is much faster in 3D because each doubling gathers 8 cells instead of 4, with a max observed speed-up of nearly 20,000$\times$ using the multilevel method.
Figure 5: Maximum error in the reconstructed velocity field generated by a 3D Hill vortex for various kernel sizes and distances from the vortex core (left). Corresponding speed-up obtained by reducing the kernel size for the multilevel algorithm (right). The mesh has dimensions $[2^{8},2^{8},2^{8}]$.
4 Confined vorticity applications
We next validate the application of the Bio-Savart conditions to the solution of a set of impulsive flow problems. The flow is truly irrotational outside the computational domain, matching the implicit assumptions required to use the Biot-Savart equations to reconstruct the velocity on the boundary.
4.1 3D flow around an accelerating disk
The flow around a disk accelerated from rest in a quiescent flow is purely potential for $t\to 0^{+}$. As a result, an analytical expression for the added mass coefficient and the pressure field can be used to validate our novel potential flow boundary condition when injected in a moment step of the Navier-Stokes equations. We place a disk of radius $R$ in an initially quiescent flow. The disk is prescribed with a constant acceleration $a$ during one non-dimensional time $t^{*}=at/U$
$U(t)=\begin{cases}at,\quad\text{if }at/U\leq 1,\\ U,\quad\text{else}.\end{cases}$ (32)
We integrate the Navier-Stokes equation during a single time-step from $t=0\to 0^{+}$ and compute the pressure force acting on the disk for various domain sizes. From potential flow theory, the added-mass coefficient for an impulsively started flow is
$Ca=\frac{F_{am}}{\rho D^{3}a}=\frac{F_{am}}{\rho D^{2}U^{2}}\left(\frac{1}{a^{*}}\right)=\frac{1}{3}$ (33)
where $Ca$ is the classical added-mass coefficient for a circular plate. We then continue the simulation until $t^{*}=3$. Once the disk reaches its peak velocity, the Reynolds number is $Re=\frac{UR}{\nu}=1000$.
We note here that the problem possesses two axes of symmetry that would allow only $1/4$ of the domain to be modelled. This can be easily accounted for in the Biot-Savart boundary conditions by adding the contributions of the images of each point. However, with our multilevel approach, they might lie on different levels, rendering the evaluation cumbersome. To avoid unnecessary complications, we model the full plate and do not use the symmetries of the problem.
Figure 6-7 shows the vorticity and pressure field at three consecutive times during the cylinder’s motion. Fig. 6 show the results with the standard reflective boundary conditions (the entire computational domain is shown), and in Fig. 7, the flow has been obtained with the Biot-Savart boundary conditions. High blockage effects are apparent for the reflective boundary conditions, with stronger shear layers and a vortex center located further downstream than the Biot-Savart results. The pressure field is even more striking. With iso-contour of the pressure sticking to the wall to satisfy the homogeneous Neumann condition set by the velocity field, see . Because the Biot-Savart boundary conditions improve this boundary condition, the pressure follows the expected potential flow solution and, more importantly, is relaxed from the homogeneous Neumann condition that reflective boundary conditions imposed.
Figure 6: Flow around an initially stationary disk accelerated in a quiescent flow at three convective times, $t^{*}\in[1,2,3]$, using the reflective boundary conditions. The entire computational domain is shown. The top half shows 10 iso-contour of the vorticity, equally spaced in the interval $\omega L/U\in\pm 20$. The bottom half shows the pressure field, with 10 iso-contour equally spaced in $p/\rho U^{2}\in\pm 2$.
Fig. 8 shows the time trace of the pressure force acting on the accelerated disk. The difference in the pressure field is found again in the difference in the forces. Here, the Biot-Savart boundary condition can exactly recover the added-mass force on the first time step, while the reflective boundary condition overestimates it. For larger $t^{*}$, the effect of the boundary condition is still observable; the reflective boundary condition significantly overestimates the pressure forces
Figure 7: Flow around an initially stationary disk accelerated in a quiescent flow at three convective times, $t^{*}\in[1,2,3]$, using the Biot-Savart boundary conditions. The entire computational domain is shown. The top half shows 10 iso-contour of the vorticity, equally spaced in the interval $\omega L/U\in\pm 20$. The bottom half shows the pressure field, with 10 iso-contour equally spaced in $p/\rho U^{2}\in\pm 2$.
Figure 8: Instantaneous pressure force acting on the disk for the two different boundary conditions. The convective time corresponds to the snapshots shown in Fig. 6-7.
[…]
Figure 9 shows the vorticity generated by this accelerating body as six convective times $t^{*}\in[1,2,3,4,5,6]$. The vortices generated are shown via the $\lambda_{2}$-criterion [21]. The initial impulse generates a strong shear layer around the plate’s edge that induces a strong vortex. Although the platform of the plate is square, this initial vortex rearrange itself into a shape resembling that of a vortex ring. After this initial vortex has traveled downstream, the plate’s corner starts shedding smaller vortices that later trigger the breakdown of the start vortex. An animation of these still frames is provided in the supplementary material.
5 Vortex-street applications
Numerous examples of external flow violate our initial assumption that the vorticity is confined near the immersed body, far from the domain’s boundary. We use two classical examples of vortex-street flows to investigate the effect of our short computational domain on flow history. First, the flow around a 2D cylinder at a Reynolds number $Re=200$, which results in the classical van Kàrmàn wake. Then, the reversed van Kàrmàn (or propulsive) wake generated by a heaving airfoil in a uniform flow at a range of Strouhal numbers.
5.1 Flow around a 2D cylinder at $Re=200$
We start our analysis with the flow behind a 2D circular cylinder of diameter $D$. The cylinder is immersed in a viscous fluid with free-stream velocity $U=(1,0)$. The viscosity is set such that the Reynolds number is $Re=\frac{UD}{\nu}=200$. We perform a domain study by varying the domain sizes from $5D\times 3D$ up to $30D\times 24D$ and, for each domain, compare the reflective and Biot-Savart boundary conditions.
Figure 10 shows the total drag force acting on the cylinder for the various domain sizes and boundary conditions. The different domain sizes and the Biot-Savart boundary conditions yield the best results with an excellent collapse of the drag forces, regardless of the domain size. Simulations performed with the Biot-Savart boundary conditions are insensitive to domain size. Reflective boundary conditions induce extremely large blockage effects for small domains and only slowly converge to the Biot-Savart boundary solution for significantly larger domains.
Additionaly, we compare the results obtained with a grid aligned with the mean flow direction to results obtained on a $7L\times 7L$ domain where the flow direction is $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$ with the Biot-Savart boundary conditions. Results are also presented in Fig. 10 and are almost indistinguishable from the results using the aligned domain.
In addition to the total drag forces acting on the cylinder, we look at the vortex shedding frequency in the wake. Here again, the Biot-Savart boundary conditions show an excellent collapse of the shedding frequency, which matches that of classical bluff body vortex shedding frequency of $St\sim 0.3$ The high blockage effect induced by the reflective boundary conditions induces fast flow around the cylinder, increasing the shedding frequency to higher values with reflective boundary conditions.
Figure 10: (left) Drag force acting on the circular cylinder at $Re=200$ for various combinations of boundary conditions and domain size. The rotated domain consists of a $4D\times 4D$ domain, with the cylinder centered at $1.5D\times 1.5D$ and a free stream velocity of $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$. (right) The frequency spectrum of the wake of the cylinder for the same combination of boundary conditions and domain size shows the classical shedding frequency at $St=0.3$.
These results suggest that, for drag wakes at least, the wake history on the immersed body is very small. However, propulsive systems use thrust wakes where wake history is expected to be higher.
5.2 Wake behind a heaving foil
We continue our analysis with thrust wakes. We immerse a rigid foil of length $L$ moving with a linear velocity $\vec{U}=(U,0)$ in a viscous fluid. In addition to its steady forward speed, the foil undergoes a pure heave motion of amplitude $h_{0}$
$y(t)=h_{0}\sin(2\pi ft)$ (34)
where $h_{0}/L=0.5$ is the non-dimensional amplitude, and the frequency is $f=St\,U/h_{0}$. For Strouhal number $St<0.5$, the airfoil generates thrust, and almost no net lift, and for $St>0.6$, the wake generated by this flapping foil is strongly asymmetric and generates a significant net lift force. We perform simulations of this system for a range of Strouhal numbers and compare the mean lift force and wake obtained on a large domain ($30L\times 20L$) with standard reflective boundary conditions to the ones obtained on two different domains of size ($6L\times 4L$) and ($12L\times 8L$) that use the Biot-Savart boundary conditions. We simulate the flow around this heaving airfoil for 100 convective times to ensure that the wake stabilizes, and we gather pressure and viscous forces for the last 20 convective times. Mean lift forces are obtained by time-averaging these forces over the past 20 convective times.
For $St<0.5$, we observe almost no difference between the different boundary conditions and domain sizes. The wake is symmetric, and the airfoil produces zero mean lift forces; see Fig. 12. For $St\geq 0.5$ we start to observe wake deflection on the large domains with the reflective boundary condition.
(a) Reflection ($30L\times 20L$)
(b) Reflection ($30L\times 20L$) - near wake
(c) Biot-Savart ($6L\times 4L$)
(d) Biot Savart ($12L\times 8L$)
Figure 11: Snapshot of the deflected wake behind an airfoil at $St=0.6$ and $Re=100$ for an amplitude to chord ratio $h_{0}/L=1.0$. The upper left panel $(a)$ shows the far wake obtained with the reflective boundary condition in a large domain, the upper right panel $(b)$ shows a zoom onto the airfoil and the near wake, the bottom left $(c)$ panel shows the entire (small) Biot-Savart domain used for the computation (the dashed line represents the domain boundaries) and the lower right panel $(d)$ shows the entire (larger) Biot-Savart domain. Vorticity is shown as 10 equally spaced isocontours in the interval $\omega L/U\pm 10$.
Fig. 11 shows the vorticity field generated by the wing at $St=0.6$. The left panel shows the complete wake obtained with the large domain of size ($30L\times 20L$). The left panel is a zoomed-in version of the same simulation, using a field of view with the size of the largest Biot-Savart domain. Finally, the bottom row shows two Biot-Savart simulations with two different domain sizes. The left panel shows the smallest domain of size ($6L\times 4L$), represented by the gray dashed line, and the right panel shows a larger domain of size ($12L\times 8L$) which extends to the whole panel.
Wake history strongly influences the near wake; the large domain correctly captures this, and a strong wake deflection is observed (the first vortex shed sets the direction; in this case, it is shed in the negative $y$ direction).
Figure 12: (left) Mean lift force acting on the airfoil at $Re=100$, for a range of Strouhal numbers for the Biot-Savart and the reflective boundary condition and (right) Poincaré map of the lift and drag coefficients for $St=0.6$
[…]
While this is a limitation of our method, at least in terms of domain size, this flow is an artifact of 2D simulations and does not occur in actual flows. Indeed, experiments on finite foils have shown that these high-lift deflected jets cannot form for finite wings, see [22, 23].
6 Conclusion
In this manuscript, we presented a novel boundary condition formulation for the incompressible Navier-Stokes equations. The methods rely on a Biot-Savart integral of the vorticity inside the domain to set the velocity boundary conditions on the domain’s boundaries. We couple these boundary conditions with an explicit projection scheme to solve the momentum and continuity equation in their primitive variables form. Leveraging the fast decaying properties of the Biot-Savart kernel, we use a multilevel algorithm to reduce the number of operations required to update the boundary nodes from $O(N)$ to $O(\log N)$ and show that the error introduced by this multilevel approach is bounded and can be estimated. We demonstrate that when a body is immersed in a domain using these boundary conditions, a coupling between the immersed body and the domain’s boundary is introduced. This coupling only results in a weakly non-linear pressure Poisson equation and a simple differed correction approach can be used to solve the coupled projection/Biot-Savart step, allowing the use of standard Poisson solver.
With different examples of flow with confined vorticity, we show that the novel boundary conditions allow for a significant reduction in the size of the computational domain without any loss in accuracy. The method is able to exactly capture added-mass force on an accelerated disk with a computational domain only extending $1/2$ diameter away from the plate.
Surprisingly, for vortex street applications, the method demonstrates its robustness and perfectly captures the drag forces and vortex shedding frequencies generated by a 2D circular cylinder at $Re=200$, even with minimal domain sizes ($5D\times 3D$). Classical reflection boundary conditions requires a domain at least 10 times bigger to generate time-varying forces indistinguishable from that of the domain with our Biot-Savart boundary conditions. Lastly, we show that the method is limited when strong wake history influences the flow in the near wake. For a high-amplitude, high-Strouhal number heaving airfoil, we show that capturing the deflected wake of the flow requires significantly larger domain (with our Biot-Savart boundary conditions) and that the computational advantage is somewhat lost.
These results show that our novel boundary conditions are very effective for confined vorticity applications, such as accelerated bodies in the initial phase ($t^{*}<10$), when the vorticity is still confined to the body, or flow with very weak wake dependency.

### v2

@@
-Using Biot-Savart to shrink Eulerian domains while maintaining or improving external flow accuracy
+Using Biot-Savart boundary conditions for unbounded external flow on Eulerian meshes
 Abstract
@@
 The Biot-Savart condition is incorporated into a general solver with Eulerian velocity-pressure fields by including the pressure-induced body vorticity during the incompressible-flow projection-step.
-We use a multilevel approach to reduce the computational cost of the evaluation of the Biot-Savart integral from $O(N)$ to $O(\log N)$ and formally show that the error associated with this approach is bounded by the total vorticity contained in the lower levels and that it scales with the inverse power of the problem’s dimension.
+We use a multilevel approach to reduce the computational cost of the evaluation of the Biot-Savart integral from $O(N)$ to $O(\log N)$ and formally show that the error associated with this approach is bounded and scales with the inverse power of the problem’s dimension.
 We use two-dimensional and three-dimensional analytic vortex problems to demonstrate this scaling is correct and select appropriate multilevel coarsening parameters accordingly.
-We show that the new method exactly captures the analytical added-mass force of a square plate accelerated from rest even when the entire domain only extends 1/2 diameter from the plate.
-The new method also predicts essentially identical time-varying drag forces for a two-dimensional circle regardless of domain size, while classical boundary conditions require a domain nearly 10 times larger to converge on the new method’s result.
-Finally, we study the sensitivity of deflected wakes to the new domain boundary conditions and show that only in minimal domains, 5 times smaller than the reference domain, the omission of the far-field wake influences the flow field and the forces on the body.
+We show that the new method exactly captures the analytical added-mass force of accelerating plates and matches experimentally measured wake development even when the entire domain only extends 1/2 diameter from the plate.
+The new method also predicts essentially identical time-varying drag forces for a two-dimensional circle regardless of domain size, while classical boundary conditions require a domain nearly 48 times larger to converge on the new method’s result.
+Finally, we study the sensitivity of 2D deflected wakes to the new domain boundary conditions and show that only in minimal domains, 25 times smaller than the reference domain, the omission of the far-field wake influences the flow field and the forces on the body.
 Doubling the domain size recovers the correct results.
@@
 Biot-Savart , Boundary Condition , Projection Method , Unbounded Domains
-PACS:
-$0$ , $1$
-MSC:
-$0$ , $1$
 journal: Journal of Computational Physics
@@
 {graphicalabstract}[Uncaptioned image]
+$a)\,\,t^{*}=2.0$
+$b)\,\,t^{*}=2.5$
+$c)\,\,t^{*}=3.0$
+$d)\,\,t^{*}=4.0$
 {highlights}
-External flow boundary condition via the Biot-Savart integral enables minimal computational domains
+Unbounded flow simulations on minimal Eulerian domains via Biot-Savart boundary conditions.
 Fast ($\log N$) multilevel evaluation of the Biot-Savart integral with controllable bounded error.
-Reduction in degrees of freedom by $O(10^{3})$ compared to reflection conditions for flow with bounded vorticity.
+Fully-couples velocity-pressure description to vorticity-condition via projection step.
+Matches analytic and experimental results while reducing required domain size by 48x.
 1 Introduction
 Computational fluid dynamics provides a numerical solution to a continuum problem in an (often) unbounded domain.
-Common practice is to approximate the continuum problem with a discrete set of equations and truncate the unbounded domain to a finite computational domain.
-While the former does not usually introduce significant errors, the latter can significantly impact the numerical solution when the finite domain is not sufficiently large.
-In the context of external flows, sufficiently large means that the decaying potential flow induced by the body and the advection/diffusion of vorticity through the domain’s boundary have no influence the forces experienced by the body [1].
-Numerical simulations of external incompressible flows enforce far-field boundary conditions on the computational domain’s exterior.
+Common practice is to approximate the continuum problem with a discrete set of equations and truncate the unbounded domain to a finite computational domain with far-field boundary conditions.
+External incompressible flows enforce far-field boundary conditions on the computational domain’s exterior.
 These consist of a known velocity normal to the boundary and a known normal pressure gradient111for a description of the boundary conditions for the incompressible Navier-Stokes equations, we refer the reader to [1].
@@
 In particular, the force acting on a body can suffer from severe blockage effects when the computational domain is not sufficiently large.
+Here, sufficiently large means that the decaying potential flow induced by the body and the advection/diffusion of vorticity through the domain’s boundary have no influence on the solution [2].
 While domain sensitivity study can determine the extent of the blockage, there is no a priori rule for estimating their magnitude, making this an expansive iterative process.
-Grid stretching away from the immersed body can efficiently maximize the domain size while maintaining a high-resolution mesh around the body, see [3, 4].
-This can limit the blockage effects on the body.
-However, this method requires correct stretching ratios to avoid excessive numerical dissipation or reflection inside the domain.
-The method is also inefficient in cases where the body undergoes substantial motion within the computational domain, for example, when tracking the motion of a kite.
-Grid stretching can be replaced by a buffer layer that acts as a numerical damper which ensures that the solution decays rapidly before it reaches the boundaries [5].
-Following a similar approach, a coarse grid can provide the boundary conditions for a relatively small, well-resolved domain; this is the basis of the multi-domain methods [1].
-The method interpolates the coarse mesh solution to the walls of the finer domain, providing artificial far-field boundary conditions.
-For flows with one [6, 7] or two unbounded directions [8], coordinate transforms or analytical potential flow solution can be applied in the unbounded directions and correctly capture solution.
-Potential flow boundary conditions are prescribed following the classical Hodge decomposition of the velocity field into an irrotational and an incompressible part
+It is also computationally inefficient to add grid cells to fill such a large domain, and grid refinement near the solid boundaries introduces well-known sensitives to the grid expansion ratios and cell aspect ratios due to numerical diffusion, dispersion, reflection, and linear system stiffness.
+Alternatively, numerical dissipation can be explicitly imposed in a buffer layer on a smaller domain that forces the solution to decay rapidly before it reaches the boundaries [3].
+Domain augmentation is a more principled approach to simulate external flow, such as using a large coarse grid to provide artificial far-field boundary conditions for a small well-resolved domain [2].
+The outer flow field can also be augmented with a vortex particle-mesh method, reducing dispersive errors [4].
+Such methods enable tracking the wake for an extended time while enabling accurate modeling of the developing boundary layer on the object’s surface and can also include immersed boundaries [5].
+However, several (user-defined) parameters are required to correctly interpolate between the two domains, and the complexity of accurately coupling the domains makes theis apporach difficult to apply to existing flow solvers.
+For flows with one [6] or two unbounded directions [7, 8], coordinate transforms and/or analytical potential flow solution [9] can be applied in the unbounded directions.
+Potential flow boundary conditions are prescribed following the classical Helmholtz decomposition of the velocity field into an irrotational and an incompressible part
 $\vec{u}=\vec{u}_{\omega}+\nabla\phi+\vec{U}_{\infty}=-\nabla^{-2}\left(\nabla\times\vec{\omega}\right)+\nabla\phi+\vec{U}_{\infty}$ (1)
 where $\vec{u}_{\omega}$, $\nabla\phi$, $\vec{U}_{\infty}$ are the incompressible component, irrotational component, and the free-stream velocity, respectively.
-Fundamental solution to discrete operators (Lattice Green’s functions or LGFs) and their asymptotic expansions can be used to approximate discrete differential equations on unbounded domains [9].
-Points close to the source are evaluated with the discrete kernel (accurate) while points far away are evaluated with the asymptotic expansion (fast).
+While these methods are elegant, they rely either on a single unbounded direction (as in [6] or formulating the problem in a coordinate system where two unbounded directions are transformed to a single one (for example, from Cartesian to Cylindrical in [8]), rendering them difficult to generalize to 3D flow problems.
+Other classes of methods solve unbounded Poisson equations by incorporating the unbounded conditions as Dirichlet boundary data, either by successive relaxed iterations of the Poisson problem with the Dirichlet data computed using the previous solution [10] or by using the James-Lackner [11, 12] algorithm that solves two nested domain Poisson problems, the inner-most with homogeneous Neumann boundary conditions and second outermost domain with inhomogeneous Neumann boundary conditions computed from a boundary convolution of the inner solution with the fundamental solution to the Poisson operator.
+To accelerate the convolution, multilevel approximation [13] or using local correction methods [14] can be used.
+Although these methods enforce unbounded boundary conditions, they require iterating a Poisson problem or solving it multiple times.
+Finally, fundamental solutions to discrete operators (Lattice Green’s functions or LGFs) and their asymptotic expansions can be used to solve the discrete Poisson equations in unbounded domains [15, 16].
+Points close to the source are evaluated with the discrete kernel (accurate), while points far away are evaluated with the asymptotic expansion (fast).
 These methods rely on the unbounded properties of the LGFs and can simulate unbounded viscous incompressible flow [17] and flows with immersed-boundaries [18].
-However, they are based on the solution of inhomogeneous, linear, constant coefficient elliptic equations and are thus incompatible with some immersed-boundary methods [3, 4].
-Other methods leverage the strength of an Eulerian-based solver for the near body and Lagrangian solvers for the far wake by coupling a near wall Eulerian solver to a Vortex particle method in the wake [12].
-They enable tracking the coarser wake for an extended time while enabling accurate modeling of the developing boundary layer on the object’s surface.
-However, several (user-defined) parameters are required to correctly interpolate the Vortex particle method solution onto the Eulerian mesh.
-Extension to three-dimensional flow is possible but will significantly increase the computational cost.
-All these methods allow for a reduction of the error introduced by truncating an infinite computational domain.
-However, their generality is often limited to specific discretization of the governing equations or introduces complex coupling procedures.
-In this manuscript, we propose a novel method for imposing external potential-flow boundary conditions on the velocity and pressure fields on the exterior of a Cartesian computational domain.
-The method retains the primitive variables formulation ($\vec{u}$ and $p$) of the covering equations and is compatible with immersed boundaries.
-Additionally, it does not introduce additional Lagrangian or Eulerian mesh.
+This allows for computational domains snug to the vorticity field, but adapting these methods to to existing Eulerian or immersed boundary solvers is difficult.
+In this manuscript, we propose a novel method for imposing external potential-flow boundary conditions of a viscous flow simulation by coupling to a vorticity-based description of the flow on the boundary via the Biot-Savart equation.
+The method retains the primitive variables formulation ($\vec{u}$ and $p$) of the governing equations, is compatible with immersed boundary methods and fast pressure projection methods, and does not require additional meshes to describe the flow.
 In the next section, we describe the Biot-Savart boundary conditions and how they are injected into a projection algorithm to solve the Navier-Stokes equations.
 Next, we demonstrate the accuracy of the velocity reconstruction and the balance between speedup and accuracy in the reconstructed field.
-We then demonstrate, with confined vorticity and vortex-street examples, that the new boundary conditions significantly outperform standard reflective boundary both in terms of computational time and solution accuracy.
-We present unexpected result that the new method continues to perform well when the fundamental assumption of potential flow external to the domain is violated, and use the unstable deflected wakes generated by fast flapping foils to study the limits of the approach.
-We close this manuscript with a discussion of the method’s applicability to various flows in science and engineering.
+With confined vorticity and vortex wake examples, we demonstrate that the new boundary conditions significantly outperform standard reflective boundary conditions on significantly smaller domains.
+We present an unexpected result: the new method continues to perform well when the fundamental assumption of potential flow external to the domain is violated, and we use the unstable deflected wakes generated by fast flapping foils to study the limits of the approach.
+We finish this manuscript by discussing the method’s applicability to various flows in science and engineering.
 2 Method
@@
 The outer domain boundary is denoted at $\partial\Omega(N^{\cal S})$, where $\cal S$ is a factor linked to the dimensions of the problem.
-In 2D we have $S=1/2$, and in 3D $S=2/3$.
+In 2D we have ${\cal S}=1/2$, and in 3D ${\cal S}=2/3$.
 The primary source of vorticity in the domain is an immersed body $\mathcal{B}$ with outer (wet) surface $\partial\mathcal{B}$ and unit normal vector $\hat{n}$, see Figure 1.
@@
 The projection method starts with an explicit estimate of the intermediate velocity field $u^{*}$
-$\vec{u}^{*}=\vec{u}\,^{t}+\int_{t}^{t+\Delta t}\frac{1}{Re}\nabla^{2}\vec{u}-\left(\vec{u}\cdot\nabla\right)\vec{u}\text{ d}t$ $\quad\forall\ \vec{x}\in\Omega(N),$ (2)
+$\vec{u}^{*}=\vec{u}\,^{t}+\int_{t}^{t+\Delta t}\frac{1}{\text{Re}}\nabla^{2}\vec{u}-\left(\vec{u}\cdot\nabla\right)\vec{u}\text{ d}t$ $\quad\forall\ \vec{x}\in\Omega(N),$ (2)
 where Re is the Reynolds number of the flow.
@@
 Body-fitted methods enforce this condition directly, while immersed boundary methods enforce it indirectly; either by reformulating the governing equations near the boundary such as [21, 22] or by modifying the discretization near $\cal{B}$ [23].
-Additionally, Eq. (2) must be supplemented with far-field boundary conditions on the domain’s outer boundaries. As discussed in the introduction, the simplest of these is a reflection condition
+Additionally, Eq. (2) must be supplemented with far-field boundary conditions on the domain’s outer boundaries. The simplest of these is a reflection condition
 $\frac{\partial\vec{u}^{*}_{s}}{\partial\hat{n}}=0,\ \vec{u}^{*}_{n}=U_{n}$ $\quad\forall\ \vec{x}\in\partial\Omega(N^{\cal S}),$ (4)
@@
 This formulation implicitly assumes the flow external to the domain is irrotational and that the flow is incompressible everywhere.
-Such a formulation is exact for the flow around accelerated bodies when the time $\frac{tU}{L}\lesssim t^{\text{crit.}}$, with $t^{\text{crit.}}$ the critical time for vortex detachment [17].
+This assumption is met when all the vorticity is confined within the computational domain such as for the flow around accelerated bodies when the time $\frac{tU}{L}\lesssim t^{\text{crit.}}$, with $t^{\text{crit.}}$ the critical time for vortex detachment [25].
 However, we will show that the results of using equation (8) in an Eulerian solver are still highly accurate for flows with wakes that extend to the domain boundary.
@@
 A naive evaluation of Eq.
-9  would require $O(N)$ operations for each boundary point, meaning $O(N^{1+\cal S})$ cost overall, overwhelming any savings from reducing the size of the domain.
-In vortex-particle methods, the integral is usually evaluated with Fast Multipole Method that allows reducing this cost to $O(\log N)$ operations.
+8 would require $O(N)$ operations for each boundary point, meaning $O(N^{1+\cal S})$ cost overall, potentially overwhelming any savings from reducing the size of the domain.
+In vortex particle-mesh methods, the integral is usually evaluated with Fast Multipole Method that allows reducing this cost to $O(\log N)$ operations.
 Here, we use a multilevel algorithm to achieve the same speed-up in our Eulerian setting.
@@
 where level $1$ is the finest domain and $l$ is the coarsest, and each domain is non-overlapping and their union fully covers $\Omega$, Fig 2.
-In this work, we define nested rectangular boxes of half-width $S^{(i)}$ centered on $\vec{x}$ and scale this width with the cell size which doubles at each level, i.e.
+In this work, we use nested rectangular boxes of half-width $S^{(i)}$ centered on $\vec{x}$, and scale this width with the cell size which doubles at each level, i.e.
 $h^{(i)}=h^{(1)}2^{i-1}$.
@@
 $f(\vec{x};\vec{\omega},\Omega)\approx\sum_{i=1}^{l}f(\vec{x};\vec{\omega}^{(i)},\mathcal{D}^{(i)}).$ (12)
-Once the multilevel field $\omega^{(i)}$ is computed, the total number of operations to evaluate equation 12 for a simulation of dimension $d=2$ or $3$ is therefore $O(l\tilde{S}^{d})=O(\log N)$ since $\tilde{S}$ is constant and $l=O(\frac{1}{d}\log N)$.
-In practise, we use the circulation over a cell face $\Gamma=\omega h^{2}$ instead of the vorticity in the Biot-Savart formulation.
+Once the multilevel field $\omega^{(i)}$ is computed, the total number of operations to evaluate Eq.
+12 for a simulation of dimension $d=2$ or $3$ is therefore $O(l\tilde{S}^{d})=O(\log N)$ since $\tilde{S}$ is constant and $l=O(\frac{1}{d}\log N)$.
+In practice, we use the circulation over a cell face $\Gamma=\omega h^{2}$ instead of the vorticity in the Biot-Savart formulation.
 The circulation on a coarser level is calculated by uniformly pooling the circulation on the level above
@@
 On uniform Cartesian meshes, $|\vec{\delta r}_{j}^{(i)}|\sim 2^{l}$, while $|\vec{r}|^{2}\sim(2^{l}\tilde{S})^{2}$.
-An upper bound for the total error in the induced velocity at a point $\vec{x}$ due to the $i$th-level evaluation of the Biot-Savart integral is therefore
+An upper bound for the total error in the induced velocity at a point $\vec{x}$ due to the multilevel evaluation of the Biot-Savart integral is therefore
 $\varepsilon(\vec{x})\approx\frac{1}{2\pi\tilde{S}^{2}}\sum_{i=2}^{l}\frac{1}{2^{i}}\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}.$ (18)
@@
 2.3 Biot-Savart & Projection Algorithm
-Substitution of the Biot-Savart boundary condition (Eq. 8) into the projection scheme results in a slightly modified intermediate velocity field
-Biot-Savart intermediate step
-$\vec{u}^{*}=\vec{u}\,^{t}+\int_{t}^{t+\Delta t}\frac{1}{Re}\nabla^{2}\vec{u}-\left(\vec{u}\cdot\nabla\right)\vec{u}\text{ d}t$ $\quad\forall\ \vec{x}\in\Omega(N),$ (20)
-$\vec{u}^{*}=f(\vec{x}_{b},\nabla\times\vec{u}^{*},\Omega)$ $\quad\forall\ \vec{x}_{b}\in\partial\Omega(N^{\cal S})$ (21)
-where the exterior boundary condition have been computed from the intermediate velocity $\vec{u}^{*}$ in the domain.
-The projection step, however, results in a coupling between the body and the domain’s boundary through the pressure field
-$\nabla\cdot\frac{\Delta t}{\rho}\nabla p=\nabla\cdot\vec{u}^{*}$ (22)
-$\vec{u}^{t+\Delta t}=\vec{u}^{*}-\frac{\Delta t}{\rho}\nabla p$ (23)
-$\vec{u}^{t+\Delta t}=f(\vec{x},\nabla\times\vec{u}^{t+\Delta t},\Omega)=\vec{u}^{*}-f(\vec{x}_{b},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega)$ (24)
-Where we have omitted the domain of definition of each equation for clarity.
-Away from the body, the source term in the Biot-Savart integral in Eq.
-24  vanishes.
-However, due to the pressure boundary condition on the immersed body ($\partial p/\partial\hat{n}=\hat{n}\cdot(\nabla^{2}\vec{u}/Re-\text{D}\vec{U}_{b}/\text{D}t)$), a thin layer of vorticity is generated by the pressure gradient term that changes the induced velocity field on the outer domain’s boundary.
-The projection steps and the Biot-Savart boundary conditions now form a fixed-point system
-Biot-Savart projection step
-$\nabla\cdot\frac{\Delta t}{\rho}\nabla p=\nabla\cdot\vec{u}^{*}-\nabla\cdot f(\vec{x}_{b},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega)$ (25)
-$\vec{u}^{t+\Delta t}=\vec{u}^{*}-\frac{\Delta t}{\rho}\nabla p$ (26)
-$\vec{u}^{t+\Delta t}=f(\vec{x},\nabla\times\vec{u}^{t+\Delta t},\Omega)=\vec{u}^{*}-f(\vec{x}_{b},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega)$ (27)
-This coupling only occurs on the domain’s boundary, and the resulting system is mildly non-linear.
-This weak non-linearity can be solved efficiently with a deferred correction approach, which explicitly updates the source term of the Poisson problem during each pressure solver iteration.
-This coupling is only apparent in immersed boundary methods if the pressure boundary conditions are correctly imposed on the immersed body during the projection step; see for example [19, 4].
-Lastly, we note that the small multilevel error discussed in the previous section means that global mass conservation is no longer guaranteed on the domain boundary.
-As such, we apply a uniform correction
-$\vec{u}=\vec{u}-\frac{\oint_{\partial\Omega}\hat{n}\cdot\vec{u}\text{ d}S}{\oint_{\partial\Omega}\text{ d}S}.$ (28)
-ensuring a globally divergence-free field is passed to the pressure projection method.
+The momentum update using the new Biot-Savart domain condition proceeds as before: forming the intermediate velocity with Eq.
+2 and projecting it to be divergence free with Eq.
+6.
+However, the new boundary condition introduces a coupling between the body and domain boundaries.
+To see this, we write the equation for the boundary velocity on $\vec{x}_{b}\in\partial\Omega(N^{\cal S})$ at the new time-step as
+$\vec{u}^{t+\Delta t}(\vec{x}_{b})=f(\vec{x_{b}},\nabla\times\vec{u}^{t+\Delta t},\Omega)=f(\vec{x_{b}},\nabla\times\vec{u}^{*},\Omega)-f\left(\vec{x_{b}},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega\right)$ (20)
+where we’ve substituted Eq.
+6.
+The last term in Eq.
+20 is non-zero on cells touching the body boundary $\partial\cal B$ because, while the gradient of a field cannot induce curl in an unbounded fluid region (×$\nabla\times\nabla\phi\equiv 0$), the pressure boundary condition on the immersed body generates a thin layer of vorticity on $\partial\cal B$.
+This contribution is unknown until after the Poisson system of Eq.
+5 is solved, but updating the normal velocity on $\partial\Omega(N^{\cal S})$ will change the source term for Eq.
+5 on cells touching the domain boundary, requiring a new pressure solution.
+This coupling between the domain and body boundaries results in a mildly non-linear fixed-point system.
+This weak non-linearity can be solved efficiently with a deferred correction approach, which explicitly updates the source term of the Poisson problem during pressure solver iteration.
+We use a Multi-Grid solver for the Poisson system, as described in [26], and updating the domain velocities and corresponding Poisson residuals at the end of each V-cycle was sufficient to ensure the final pressure and velocity field satisfy both the momentum and divergence-free conditions without increasing the number of solver iterations.
+Note that this coupling is only apparent in immersed boundary methods if the pressure boundary conditions are correctly imposed on the immersed body during the projection step; see for example [27, 20].
+This approach differs from the one in [10] for unbounded Poisson problems with embedded Dirichlet boundaries as we directly adjust the residual of the (iterative) Poisson solver and we do not iterate over the coupled Poisson/interior boundary problem.
+It is also different to the James-Lackner algorithm for the same reason [13].
+Lastly, we note that the small multilevel error discussed in the previous section means that global mass conservation is no longer guaranteed to machine precision on the domain boundary.
+As such, we apply a uniform correction to the velocity computed with Eq.
+20
+$\vec{u}(\vec{x}_{b})=\vec{u}(\vec{x}_{b})-\frac{\oint_{\partial\Omega}\hat{n}\cdot\vec{u}\text{ d}S}{\oint_{\partial\Omega}\text{ d}S}.$ (21)
+ensuring a globally divergence-free field is passed to the pressure projection method on each iteration.
 3 Multilevel reconstruction validation
-We start by validating the multilevel reconstruction algorithm presented above.
-We selected to cases such respect our assumption of vorticity being confined far from the domain’s boundary.
-We demonstrate the 2D and 3D reconstruction error of the multilevel and the speedup generated by selecting the correct kernel radius from the above analysis.
+We validate the multilevel reconstruction algorithm presented above using 2D and 3D analytic vorticity fields.
+We also study the multilevel reconstruction error and the speedup generated by selecting the correct kernel radius from the above analysis.
 Our Biot-Savart boundary conditions are implemented inside WaterLily.jl [^2] https://github.com/weymouth/WaterLily.jl, a fast, immersed-boundary flow solver based on the Boundary data immersion method [22] that can execute both on CPU and GPU [28].
 To leverage the GPU capability of the flow solver and fully benefit from our novel boundary conditions, we perform all computations presented herein on an NVIDIA-A6000 GPU (48GB memory).
-All the codes used for this study and the instructions to reproduce these results are freely available at https://github.com/weymouth/BiotSavartBCs.jl.
-3.1 2D validation: Lamb vortex
-The velocity field generated by the Lamb-Chaplygin dipole is obtained from the scalar stream function $\vec{u}=-\vec{e}_{z}\times\nabla\psi$ given by
+The complete code and instructions needed to reproduce these results are freely available at https://github.com/weymouth/BiotSavartBCs.jl.
+3.1 2D validation: Lamb Chaplygin vortex-dipole
+The velocity field generated by the Lamb-Chaplygin dipole is defined by the scalar stream function
 $\psi(r)=\begin{cases}\frac{-2UJ_{1}(\beta r)}{\beta J_{0}(\beta R)},\quad\text{for}\quad r<R,\\ U\left(\frac{R^{2}}{r}-r\right),\quad\text{for}\quad r\geq R,\end{cases}$ (22)
-where $J_{0}$ and $J_{1}$ are the zeroth and first Bessel functions of the first kind, respectively.
-$\beta$ is a parameter with the value $\beta=1.2197\pi/R$.
-The core of the vortex is defined by the isoline $\psi(r=R)=0$, that delimits the inside and outside of the vortex.
-(a) $\log(\text{size})=1$
-(b) $\log(\text{size})=3$
-(c) $\log(\text{size})=5$
-(d) $\log(\text{size})=8$
-Figure 3: $L_{2}$-norm of the error in reconstructed velocity field outside the core of the 2D Lamb-Chaplygin vortex for varying (multi-level) kernel size. Error bars are shown in log-scale.
+where $J_{0}$ and $J_{1}$ are the zeroth and first Bessel functions of the first kind, respectively, and $\beta$ is a parameter with the value $\beta=1.2197\pi/R$.
+The 2D velocity field is given by $\vec{u}=-\vec{e}_{z}\times\nabla\psi$, and the core of the vortex is defined by the isoline $\psi(r=R)=0$.
+$10^{-1}$
+$10^{-2}$
+$10^{-3}$
+$10^{-4}$
+$10^{-5}$
+$10^{-6}$
+$a)$ $S=2$
+$c)$ $S=8$
+$d)$ $S=16$
+$f)$ $S=128$
+$b)$ $S=4$
+$e)$ $S=32$
+Figure 3: $L_{2}$-norm of the error in reconstructed velocity field outside the core of the 2D Lamb-Chaplygin vortex for varying (multi-level) kernel size.
 We reconstruct the velocity field outside the vortex core using the internal vorticity and our multilevel Biot-Savart method and compare that to the analytic result.
@@
 3.
-Setting the refine size $\tilde{S}$ to the domain size $\cal D$ eliminates the multi-level-error, but the result is not exact since the analytic vorticity distribution is sampled on a finite grid of size $\Delta x=0.05R$.
+Setting the refine size $S$ to the domain size $\cal D$ eliminates the multi-level-error, but the result is not exact since the analytic vorticity distribution is sampled on a finite grid of size $\Delta x=0.05R$.
 This error depends primarily on the distance from the vortex, Fig.
-3(d).
-Progressively reducing the refinement size $\tilde{S}$ introduces the multilevel error quantified in the previous section, which is only weakly dependent on the distance to the vortex.
+3(f).
+Progressively reducing the refinement size $S$ introduces the multilevel error quantified in the previous section, which is only weakly dependent on the distance to the vortex.
 This is further quantified in Fig.
 4(a), which shows the maximum error in the reconstructed velocity field at a distance $d$ from the vortex from the Lamb-Chaplygin vortex.
-The size $\tilde{S}$ is shown to dominate the reconstruction error, with fairly uniform total error levels once $\tilde{S}\leq 2^{3}=8$.
+The size $S$ is shown to dominate the reconstruction error, with fairly uniform total error levels once $S\leq 2^{3}=8$.
 Fig.
-4(a) reports the measured computational speed-up achieved by the multilevel method relative to setting $\tilde{S}=\cal D$, demonstrating up to 300$\times$ acceleration.
+4(b) reports the measured computational speed-up achieved by the multilevel method relative to setting $S=\Omega$, demonstrating up to 300$\times$ acceleration.
 Figure 4: Maximum error in the reconstructed velocity field generated by a 2D Lamb-Chaplygin vortex dipole for various kernel sizes and distances from the vortex core (left).
@@
 with the corresponding errors and speed-up shown in Figure 5.
-Note that the analytic vorticity of the hill vortex is discontinuous, increasing the truncation error near the vortex compared to the smooth Lamp-vortex, a problem which would not be observed with any real three-dimensional viscous flow.
+Note that the analytic vorticity of the hill vortex is discontinuous, increasing the truncation error near the vortex compared to the smooth 2D Lamb-dipole.
 As argued in Sec.
-2.3 the multilevel error decreases extremely quickly as $\tilde{S}$ is increased for three-dimensional flows.
+2.3 the multilevel error decreases extremely quickly as $S$ is increased for three-dimensional flows.
 This is due to the scaling of the Biot-Savart kernel itself.
@@
 4 Confined vorticity applications
-We next validate the application of the Bio-Savart conditions to the solution of a set of impulsive flow problems.
+Next, we validate the application of the Biot-Savart conditions to a set of impulsive flow problems.
 The flow is truly irrotational outside the computational domain, matching the implicit assumptions required to use the Biot-Savart equations to reconstruct the velocity on the boundary.
-4.1 3D flow around an accelerating disk
-The flow around a disk accelerated from rest in a quiescent flow is purely potential for $t\to 0^{+}$.
-As a result, an analytical expression for the added mass coefficient and the pressure field can be used to validate our novel potential flow boundary condition when injected in a moment step of the Navier-Stokes equations.
-We place a disk of radius $R$ in an initially quiescent flow.
-The disk is prescribed with a constant acceleration $a$ during one non-dimensional time $t^{*}=at/U$
+4.1 2D and 3D flow around an accelerating disk
+The flow around a 2D or 3D disk accelerated from rest in a quiescent flow is purely potential for $t\to 0^{+}$.
+As a result, an analytical expression for the added mass coefficient can be used to validate our novel potential flow boundary condition when injected in a moment step of the Navier-Stokes equations.
+The disk of radius $R$ and diameter $D=2R$ is given a constant acceleration $a$ for the initial non-dimensional time $t^{*}=at/U\leq 1$ and is given a constant speed after this, ie
 $U(t)=\begin{cases}at,\quad\text{if }at/U\leq 1,\\ U,\quad\text{else}.\end{cases}$ (25)
-We integrate the Navier-Stokes equation during a single time-step from $t=0\to 0^{+}$ and compute the pressure force acting on the disk for various domain sizes.
+We perform 2D and 3D simulations of the flow around a plate/circular disk immersed in a domain of dimension $1.5D\times 1.5D(\times 1.5D)$.
 From potential flow theory, the added-mass coefficient for an impulsively started flow is
-$Ca=\frac{F_{am}}{\rho D^{3}a}=\frac{F_{am}}{\rho D^{2}U^{2}}\left(\frac{1}{a^{*}}\right)=\frac{1}{3}$ (33)
-where $Ca$ is the classical added-mass coefficient for a circular plate.
-We then continue the simulation until $t^{*}=3$.
-Once the disk reaches its peak velocity, the Reynolds number is $Re=\frac{UR}{\nu}=1000$.
-We note here that the problem possesses two axes of symmetry that would allow only $1/4$ of the domain to be modelled.
-This can be easily accounted for in the Biot-Savart boundary conditions by adding the contributions of the images of each point.
-However, with our multilevel approach, they might lie on different levels, rendering the evaluation cumbersome.
-To avoid unnecessary complications, we model the full plate and do not use the symmetries of the problem.
-Figure 6-7 shows the vorticity and pressure field at three consecutive times during the cylinder’s motion.
-Fig.
-6 show the results with the standard reflective boundary conditions (the entire computational domain is shown), and in Fig.
-7, the flow has been obtained with the Biot-Savart boundary conditions.
-High blockage effects are apparent for the reflective boundary conditions, with stronger shear layers and a vortex center located further downstream than the Biot-Savart results.
-The pressure field is even more striking.
-With iso-contour of the pressure sticking to the wall to satisfy the homogeneous Neumann condition set by the velocity field, see .
-Because the Biot-Savart boundary conditions improve this boundary condition, the pressure follows the expected potential flow solution and, more importantly, is relaxed from the homogeneous Neumann condition that reflective boundary conditions imposed.
-Figure 6: Flow around an initially stationary disk accelerated in a quiescent flow at three convective times, $t^{*}\in[1,2,3]$, using the reflective boundary conditions.
-The entire computational domain is shown.
+$Ca=\frac{F_{am}}{\rho D^{3}a}=\frac{F_{am}}{\rho D^{2}U^{2}}\left(\frac{1}{a^{*}}\right)$ (26)
+where $Ca$ is the classical added-mass coefficient. We simulate the 2D or 3D flow until $t^{*}=3$. Once the disk reaches its peak velocity, the Reynolds number is $\text{Re}=\frac{UR}{\nu}=1000$.
+We note here that the 2D and 3D problems possess axes of symmetry that would allow only a portion of the domain to be modeled.
+This can be easily accounted for in the Biot-Savart boundary conditions by adding the contributions of the images of each vortex cell, see A.
+However, to avoid unnecessary complications in this manuscript, we model the full plate and do not use the problem’s symmetries.
+Figure 6 shows the vorticity and pressure field at three consecutive times during the disk’s motion on the midplane of the 3D circular disk.
+The top row shows the results with the standard reflective boundary conditions (the entire computational domain is shown), and the bottom row shows the flow obtained with the Biot-Savart boundary conditions.
+High blockage effects are apparent for the reflective boundary conditions, which generate a stronger shear layer that feeds a stronger vortex located further downstream than the Biot-Savart results.
+The error in the pressure field due to the homogeneous Neumann condition placed close to the immersed body generates pressure peaks twice as large as with the Biot-Savart boundary conditions.
+Figure 6: Flow around an initially stationary disk accelerated in a quiescent flow at three convective times, $t^{*}\in[1,2,3]$, using the reflective boundary conditions (top) and the Biot-Savart boundary condition (bottom).
+A slice through the middle of the entire computational domain is shown.
 The top half shows 10 iso-contour of the vorticity, equally spaced in the interval $\omega L/U\in\pm 20$.
@@
 Fig.
-8 shows the time trace of the pressure force acting on the accelerated disk.
+7 shows the time trace of the pressure force acting on the accelerated disk, as well as the equivalent 2D simulation on an accelerating flat plate for comparison.
 The difference in the pressure field is found again in the difference in the forces.
 Here, the Biot-Savart boundary condition can exactly recover the added-mass force on the first time step, while the reflective boundary condition overestimates it.
-For larger $t^{*}$, the effect of the boundary condition is still observable; the reflective boundary condition significantly overestimates the pressure forces
-Figure 7: Flow around an initially stationary disk accelerated in a quiescent flow at three convective times, $t^{*}\in[1,2,3]$, using the Biot-Savart boundary conditions.
-The entire computational domain is shown.
-The top half shows 10 iso-contour of the vorticity, equally spaced in the interval $\omega L/U\in\pm 20$.
-The bottom half shows the pressure field, with 10 iso-contour equally spaced in $p/\rho U^{2}\in\pm 2$.
-Figure 8: Instantaneous pressure force acting on the disk for the two different boundary conditions. The convective time corresponds to the snapshots shown in Fig. 6-7.
+For larger $t^{*}$, the effect of the boundary condition is still observable; the reflective boundary condition significantly overestimates the pressure forces for all $t$.
+Figure 7: Instantaneous pressure force acting on the disk for the two different boundary conditions. The convective time corresponds to the snapshots shown in Fig. 6
 […]
 […]
+We qualitatively compare the shape of the vortex ring forming behind our accelerating square plate with the flow visualization of the same cases in [29], who used dye coloring to visualize the wake forming behind the accelerating plate.
+Results are presented in Figure 9-8 for four instants during and after the acceleration phase.
+$a)\,\,t^{*}=2.0$
+$b)\,\,t^{*}=2.5$
+$c)\,\,t^{*}=3.0$
+$d)\,\,t^{*}=4.0$
 […]
-Figure 9 shows the vorticity generated by this accelerating body as six convective times $t^{*}\in[1,2,3,4,5,6]$.
-The vortices generated are shown via the $\lambda_{2}$-criterion [21].
-The initial impulse generates a strong shear layer around the plate’s edge that induces a strong vortex.
-Although the platform of the plate is square, this initial vortex rearrange itself into a shape resembling that of a vortex ring.
-After this initial vortex has traveled downstream, the plate’s corner starts shedding smaller vortices that later trigger the breakdown of the start vortex.
-An animation of these still frames is provided in the supplementary material.
+An excellent agreement between the simulations and the experiments is found for this qualitative verification, both in the oblique and downstream view from the plate.
+The initial vortex roll-up starts to break down due to the high-strength vortices generated by the plate’s edges.
+The initial roll-up is extremely similar between both visualization techniques, and the shape of the vortex ring is extremely well captured by the simulations despite the minimal domain size.
+$a)\,\,t^{*}=2.0$
+$b)\,\,t^{*}=2.5$
+$c)\,\,t^{*}=3.0$
+$d)\,\,t^{*}=4.0$
+[…]
 5 Vortex-street applications
-Numerous examples of external flow violate our initial assumption that the vorticity is confined near the immersed body, far from the domain’s boundary.
-We use two classical examples of vortex-street flows to investigate the effect of our short computational domain on flow history.
-First, the flow around a 2D cylinder at a Reynolds number $Re=200$, which results in the classical van Kàrmàn wake.
-Then, the reversed van Kàrmàn (or propulsive) wake generated by a heaving airfoil in a uniform flow at a range of Strouhal numbers.
-5.1 Flow around a 2D cylinder at $Re=200$
+Next, we study the accuracy of the new Biot-Savart conditions on flow cases with vortex wakes that are truncated by the domain boundary, introducing an additional modelling error.
+We use two classical examples of vortex-street flows: (1) the flow around a 2D cylinder at a Reynolds number $\text{Re}=200$, which results in the classical van Kàrmàn wake, and (2) the reversed van Kàrmàn (or propulsive) wake generated by a heaving airfoil in a uniform flow at a range of Strouhal numbers.
+5.1 Flow around a 2D cylinder at $\text{Re}=200$
 We start our analysis with the flow behind a 2D circular cylinder of diameter $D$.
 The cylinder is immersed in a viscous fluid with free-stream velocity $U=(1,0)$.
-The viscosity is set such that the Reynolds number is $Re=\frac{UD}{\nu}=200$.
+The viscosity is set such that the Reynolds number is $\text{Re}=\frac{UD}{\nu}=200$.
 We perform a domain study by varying the domain sizes from $5D\times 3D$ up to $30D\times 24D$ and, for each domain, compare the reflective and Biot-Savart boundary conditions.
 Figure 10 shows the total drag force acting on the cylinder for the various domain sizes and boundary conditions.
-The different domain sizes and the Biot-Savart boundary conditions yield the best results with an excellent collapse of the drag forces, regardless of the domain size.
-Simulations performed with the Biot-Savart boundary conditions are insensitive to domain size.
-Reflective boundary conditions induce extremely large blockage effects for small domains and only slowly converge to the Biot-Savart boundary solution for significantly larger domains.
-Additionaly, we compare the results obtained with a grid aligned with the mean flow direction to results obtained on a $7L\times 7L$ domain where the flow direction is $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$ with the Biot-Savart boundary conditions.
+The Biot-Savart boundary conditions yield the best results with an excellent collapse of the drag forcesregardless of the domain size, including the tiny $5Dx3D$ domain which captures no more than two wake vortices at any time.
+Reflective boundary conditions induce extremely large blockage effects for small domains and only slowly converge to the Biot-Savart boundary solution with a domain 48x smaller.
+Additionally, we compare the results obtained with a grid aligned with the mean flow direction to results obtained on a $7D\times 7D$ domain where the flow direction is $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$ with the Biot-Savart boundary conditions.
 Results are also presented in Fig.
@@
 In addition to the total drag forces acting on the cylinder, we look at the vortex shedding frequency in the wake.
-Here again, the Biot-Savart boundary conditions show an excellent collapse of the shedding frequency, which matches that of classical bluff body vortex shedding frequency of $St\sim 0.3$ The high blockage effect induced by the reflective boundary conditions induces fast flow around the cylinder, increasing the shedding frequency to higher values with reflective boundary conditions.
-Figure 10: (left) Drag force acting on the circular cylinder at $Re=200$ for various combinations of boundary conditions and domain size.
+Here again, the Biot-Savart boundary conditions show an excellent collapse of the shedding frequency, which matches that of classical bluff body vortex shedding frequency of $St\sim 0.3$.
+The high blockage effect induced by the reflective boundary conditions induces fast flow around the cylinder, adding a non-physical domain-size dependance on the shedding frequency.
+Figure 10: (left) Drag force acting on the circular cylinder at $\text{Re}=200$ for various combinations of boundary conditions and domain size.
 The rotated domain consists of a $4D\times 4D$ domain, with the cylinder centered at $1.5D\times 1.5D$ and a free stream velocity of $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$.
 (right) The frequency spectrum of the wake of the cylinder for the same combination of boundary conditions and domain size shows the classical shedding frequency at $St=0.3$.
-These results suggest that, for drag wakes at least, the wake history on the immersed body is very small. However, propulsive systems use thrust wakes where wake history is expected to be higher.
 5.2 Wake behind a heaving foil
-We continue our analysis with thrust wakes.
+The previous results suggest that the wake history on the immersed body is very small for drag wakes.
+We next continue our analysis with thrust wakes, where the influence of wake history is expected to be stronger.
 We immerse a rigid foil of length $L$ moving with a linear velocity $\vec{U}=(U,0)$ in a viscous fluid.
@@
 $y(t)=h_{0}\sin(2\pi ft)$ (27)
-where $h_{0}/L=0.5$ is the non-dimensional amplitude, and the frequency is $f=St\,U/h_{0}$.
-For Strouhal number $St<0.5$, the airfoil generates thrust, and almost no net lift, and for $St>0.6$, the wake generated by this flapping foil is strongly asymmetric and generates a significant net lift force.
+where $h_{0}/L=0.5$ is the non-dimensional amplitude, and the frequency is $f=\text{St}\,U/h_{0}$.
+For Strouhal number $\text{St}<0.5$, the airfoil generates thrust, and almost no net lift, and for $\text{St}>0.6$, the wake generated by this flapping foil is strongly asymmetric and generates a significant net lift force.
 We perform simulations of this system for a range of Strouhal numbers and compare the mean lift force and wake obtained on a large domain ($30L\times 20L$) with standard reflective boundary conditions to the ones obtained on two different domains of size ($6L\times 4L$) and ($12L\times 8L$) that use the Biot-Savart boundary conditions.
@@
 Mean lift forces are obtained by time-averaging these forces over the past 20 convective times.
-For $St<0.5$, we observe almost no difference between the different boundary conditions and domain sizes.
-The wake is symmetric, and the airfoil produces zero mean lift forces; see Fig.
-12.
-For $St\geq 0.5$ we start to observe wake deflection on the large domains with the reflective boundary condition.
 (a) Reflection ($30L\times 20L$)
@@
 (d) Biot Savart ($12L\times 8L$)
-Figure 11: Snapshot of the deflected wake behind an airfoil at $St=0.6$ and $Re=100$ for an amplitude to chord ratio $h_{0}/L=1.0$.
-The upper left panel $(a)$ shows the far wake obtained with the reflective boundary condition in a large domain, the upper right panel $(b)$ shows a zoom onto the airfoil and the near wake, the bottom left $(c)$ panel shows the entire (small) Biot-Savart domain used for the computation (the dashed line represents the domain boundaries) and the lower right panel $(d)$ shows the entire (larger) Biot-Savart domain.
+Figure 11: Snapshot of the deflected wake behind an airfoil at $\text{St}=0.6$ and $\text{Re}=100$ for an amplitude to chord ratio $h_{0}/L=1.0$.
+The upper left panel $(a)$ shows the far wake obtained with the reflective boundary condition in a large domain, the upper right panel $(b)$ shows a zoom onto the airfoil and the near wake, the bottom left $(c)$ panel shows the entire (small) Biot-Savart domain used for the computation and the lower right panel $(d)$ shows the entire (larger) Biot-Savart domain.
 Vorticity is shown as 10 equally spaced isocontours in the interval $\omega L/U\pm 10$.
-Fig.
-11 shows the vorticity field generated by the wing at $St=0.6$.
-The left panel shows the complete wake obtained with the large domain of size ($30L\times 20L$).
-The left panel is a zoomed-in version of the same simulation, using a field of view with the size of the largest Biot-Savart domain.
-Finally, the bottom row shows two Biot-Savart simulations with two different domain sizes.
-The left panel shows the smallest domain of size ($6L\times 4L$), represented by the gray dashed line, and the right panel shows a larger domain of size ($12L\times 8L$) which extends to the whole panel.
-Wake history strongly influences the near wake; the large domain correctly captures this, and a strong wake deflection is observed (the first vortex shed sets the direction; in this case, it is shed in the negative $y$ direction).
-Figure 12: (left) Mean lift force acting on the airfoil at $Re=100$, for a range of Strouhal numbers for the Biot-Savart and the reflective boundary condition and (right) Poincaré map of the lift and drag coefficients for $St=0.6$
+Figure 12: (left) Mean lift force acting on the airfoil at $\text{Re}=100$, for a range of Strouhal numbers for the Biot-Savart and the reflective boundary condition and (right) Poincaré map of the lift and drag coefficients for $St=0.6$
 […]
-While this is a limitation of our method, at least in terms of domain size, this flow is an artifact of 2D simulations and does not occur in actual flows.
+We expect that in any case where wake history effects are highly dominant, the new Biot-Savart boundary condition will require a domain large enough for this wake to develop “stable” dynamics before exiting the domain.
+However, we should note that the case of 2D deflected thrust wakes is somewhat pathological, depending on long-range vortex interactions only possible in 2D flows.
 Indeed, experiments on finite foils have shown that these high-lift deflected jets cannot form for finite wings, see [30, 31].
 6 Conclusion
-In this manuscript, we presented a novel boundary condition formulation for the incompressible Navier-Stokes equations.
-The methods rely on a Biot-Savart integral of the vorticity inside the domain to set the velocity boundary conditions on the domain’s boundaries.
+In this manuscript, we presented a novel boundary condition formulation for external incompressible flow simulations.
+The method relies on a Biot-Savart integral of the vorticity inside the domain to set the velocity boundary conditions on the domain boundaries.
 We couple these boundary conditions with an explicit projection scheme to solve the momentum and continuity equation in their primitive variables form.
@@
 The method is able to exactly capture added-mass force on an accelerated disk with a computational domain only extending $1/2$ diameter away from the plate.
-Surprisingly, for vortex street applications, the method demonstrates its robustness and perfectly captures the drag forces and vortex shedding frequencies generated by a 2D circular cylinder at $Re=200$, even with minimal domain sizes ($5D\times 3D$).
-Classical reflection boundary conditions requires a domain at least 10 times bigger to generate time-varying forces indistinguishable from that of the domain with our Biot-Savart boundary conditions.
+Surprisingly, for vortex street applications, the method demonstrates its robustness and perfectly captures the drag forces and vortex shedding frequencies generated by a 2D circular cylinder at $\text{Re}=200$, even with minimal domain sizes ($5D\times 3D$).
+Classical reflection boundary conditions require a domain 48 times bigger to generate time-varying forces indistinguishable from that of the domain with our Biot-Savart boundary conditions.
 Lastly, we show that the method is limited when strong wake history influences the flow in the near wake.
-For a high-amplitude, high-Strouhal number heaving airfoil, we show that capturing the deflected wake of the flow requires significantly larger domain (with our Biot-Savart boundary conditions) and that the computational advantage is somewhat lost.
-These results show that our novel boundary conditions are very effective for confined vorticity applications, such as accelerated bodies in the initial phase ($t^{*}<10$), when the vorticity is still confined to the body, or flow with very weak wake dependency.
+For a high-amplitude, high-Strouhal number heaving airfoil, we show that capturing the deflected wake of the flow requires significantly larger domain, implying that the computational advantage of the new boundary condition is reduced.
+The generality and accuracy of the new approach, even when simulating flows with vortex wakes that are truncated by the domain boundaries, opens a huge range of potential applications.
+In addition to the welcomed speed-up enabled by using frankly tiny computational domains, the continued accuracy of the simulations challenges the notion of the importance of far-field wake dynamics on the near-body flow, providing compelling evidence to the contrary.
+Appendix A Accounting for symmetries in the Biot-Savart integral
+We briefly describe how symmetries can be accounted for in the multilevel evaluation of the Biot-Savart integral Eq.
+8.
+We focus on a description of the 2D case; the extension to the 3D is trivial.
+For a flow with symmetries, like the 2D flow around a flat plate, see Figure 13, whose axis of symmetries is represented by the dash-dotted line, the computational domain $\Omega$ only models the upper part of the plate.
+The domain $\Omega^{\dagger}$ is the image of the computational domain $\Omega$ about the symmetry axis.
+Evaluation of the induced velocity at a point on the boundary of the computational domain $\vec{x}\in\partial\Omega$ (apart from points on the symmetry plane which have the reflection condition applied to them) on the $i$th-level is given by
+${\vec{u}}({\vec{x}})=\int_{\mathcal{R}^{(i)}}K_{n}({\vec{x}}-\vec{y})\times\hat{e}_{z}\omega({\vec{y}})\text{ d}\vec{y}+\int_{{\mathcal{R}^{(i)}}^{\dagger}}K_{n}({\vec{x}}-\vec{y}^{\dagger})\times\hat{e}_{z}\omega^{\dagger}({\vec{y}}^{\dagger})\text{ d}\vec{y}^{\dagger}$ (28)
+where $\mathcal{R}^{(i)}\equiv\mathcal{D}^{(i)}\cap\Omega$ and ${\mathcal{R}^{(i)}}^{\dagger}\equiv\mathcal{D}^{(i)}\cap\Omega^{\dagger}$.
+Additionally, $\overline{\mathcal{R}^{(i)}}^{\dagger}\subset\mathcal{R}^{(i)}$ is the images of ${\mathcal{R}^{(i)}}^{\dagger}$ in $\Omega$ and ${\vec{y}}^{\dagger}$ is the images of $\vec{y}$ in $\Omega^{\dagger}$, respectively.
+Points inside ${\mathcal{R}^{(i)}}$ (blue region) are evaluated on the computational domain $\Omega$.
+For points in ${\mathcal{R}^{(i)}}^{\dagger}$ (orange region), the image vorticity is computed by applying the symmetric condition to the vorticity in $\overline{\mathcal{R}^{(i)}}^{\dagger}$ (pink region)
+$\omega^{\dagger}(\vec{y}^{\dagger})=-\omega(\vec{y}),\quad\quad\vec{y}\in\overline{\mathcal{R}^{(i)}}^{\dagger},$ (29)
+for 2D flows; for 3D flows, computing the image vorticity vector requires computing its reflection about the symmetry plane.
+Substituting in Eq.
+28 yields an integral only over the region $\mathcal{R}^{(i)}$
+${\vec{u}}({\vec{x}})=\int_{\mathcal{R}^{(i)}}\left(K_{n}({\vec{x}}-\vec{y})-K_{n}({\vec{x}}-\vec{y}^{\dagger})\right)\times\hat{e}_{z}\omega({\vec{y}})\text{ d}\vec{y},$ (30)
+where the second term in the parenthesis only appears if $\mathcal{D}^{(i)}$ overlaps with the domain’s image $\Omega^{\dagger}$.
+A similar expression is found for 3D flows.
+As Eq.
+30  solely relies on the vorticity in the computational domain $\Omega$, no additional computations are required, and accounting for the image’s induced velocity simply requires accounting for its influence through the kernel $K_{n}$.
+$\mathcal{D}^{(i)}$
+$\Omega$
+$\vec{y}$
+$\vec{y}\,^{\dagger}$
+$\vec{x}$
+${S}\,^{(i)}$
+$\mathcal{R}^{(i)}$
+${\mathcal{R}^{(i)}}^{\dagger}$
+$\overline{\mathcal{R}^{(i)}}^{\dagger}$
+${S}\,^{(i)}$
+$\Omega^{\dagger}$
+$\partial\Omega$
+Figure 13: Schematic of the approach used to account for symmetries in the multilevel evaluation of the Biot-Savart integral.
+The induced velocity at a point $\vec{x}$ includes the contribution from $\vec{y}$ and its image point $\vec{y}^{\dagger}$.
+The computational domain $\Omega$ and its image $\Omega^{\dagger}$ intersect with a multilevel domain $\mathcal{D}^{(i)}$ of half-width $S^{(i)}$ producing the regions $\mathcal{R}^{(i)}$, ${\mathcal{R}^{(i)}}^{\dagger}$ and $\overline{\mathcal{R}^{(i)}}^{\dagger}$.

### v3

@@
 Abstract
-We introduce a novel boundary condition formulated using a Biot-Savart vorticity integral inside an Eulerian fluid domain that maintains high-accuracy results even when the domain boundary is within a body-length of immersed solid boundaries.
-The Biot-Savart condition is incorporated into a general solver with Eulerian velocity-pressure fields by including the pressure-induced body vorticity during the incompressible-flow projection-step.
-We use a multilevel approach to reduce the computational cost of the evaluation of the Biot-Savart integral from $O(N)$ to $O(\log N)$ and formally show that the error associated with this approach is bounded and scales with the inverse power of the problem’s dimension.
-We use two-dimensional and three-dimensional analytic vortex problems to demonstrate this scaling is correct and select appropriate multilevel coarsening parameters accordingly.
-We show that the new method exactly captures the analytical added-mass force of accelerating plates and matches experimentally measured wake development even when the entire domain only extends 1/2 diameter from the plate.
-The new method also predicts essentially identical time-varying drag forces for a two-dimensional circle regardless of domain size, while classical boundary conditions require a domain nearly 48 times larger to converge on the new method’s result.
-Finally, we study the sensitivity of 2D deflected wakes to the new domain boundary conditions and show that only in minimal domains, 25 times smaller than the reference domain, the omission of the far-field wake influences the flow field and the forces on the body.
-Doubling the domain size recovers the correct results.
+[…]
 keywords:
@@
 […]
-[inst1]organization=Ship Hydromechanics,addressline=Faculty of Mechanical Engineering (ME),
+[inst1]addressline=Faculty of Mechanical Engineering (ME),
 Delft University of Technology, city=Delft, country=Netherlands
@@
 {highlights}
-Unbounded flow simulations on minimal Eulerian domains via Biot-Savart boundary conditions.
-Fast ($\log N$) multilevel evaluation of the Biot-Savart integral with controllable bounded error.
-Fully-couples velocity-pressure description to vorticity-condition via projection step.
-Matches analytic and experimental results while reducing required domain size by 48x.
+Biot-Savart boundary conditions produce accurate unbounded flow simulations on Eulerian domains minimally larger than the immersed solid geometries.
+Iterative partitioned solution of the non-locally coupled immersed surfaces and domain surfaces in the projection step requires minimal adjustment to classic velocity-pressure incompressible Navier-Stokes solvers.
+Simplified oct-tree clustering for the Fast Multi-level Method ($\ell$M) integral evaluation has bounded error, optimal $O(N)$ operations, and efficient parallel execution on GPUs.
 1 Introduction
-Computational fluid dynamics provides a numerical solution to a continuum problem in an (often) unbounded domain.
-Common practice is to approximate the continuum problem with a discrete set of equations and truncate the unbounded domain to a finite computational domain with far-field boundary conditions.
-External incompressible flows enforce far-field boundary conditions on the computational domain’s exterior.
-These consist of a known velocity normal to the boundary and a known normal pressure gradient111for a description of the boundary conditions for the incompressible Navier-Stokes equations, we refer the reader to [1].
-These can lead to significant errors.
-In particular, the force acting on a body can suffer from severe blockage effects when the computational domain is not sufficiently large.
-Here, sufficiently large means that the decaying potential flow induced by the body and the advection/diffusion of vorticity through the domain’s boundary have no influence on the solution [2].
-While domain sensitivity study can determine the extent of the blockage, there is no a priori rule for estimating their magnitude, making this an expansive iterative process.
-It is also computationally inefficient to add grid cells to fill such a large domain, and grid refinement near the solid boundaries introduces well-known sensitives to the grid expansion ratios and cell aspect ratios due to numerical diffusion, dispersion, reflection, and linear system stiffness.
-Alternatively, numerical dissipation can be explicitly imposed in a buffer layer on a smaller domain that forces the solution to decay rapidly before it reaches the boundaries [3].
-Domain augmentation is a more principled approach to simulate external flow, such as using a large coarse grid to provide artificial far-field boundary conditions for a small well-resolved domain [2].
+When simulating external viscous flow problems with an Eulerian method using velocity and pressure as the primary variables, the ideally unbounded domain must be truncated to a finite size and artificial domain boundary conditions must be imposed on the computational domain’s exterior.
+These typically consist of a known velocity normal to the boundary and a known normal pressure gradient [1].
+These can lead to significant blockage errors when the computational domain is not sufficiently large, restricting the decay of the potential flow induced by the body, or advecting/diffusing vorticity through the domain’s boundary [2].
+While domain sensitivity study can determine the extent of the blockage, there is no a priori rule for estimating their magnitude, and it is computationally inefficient to add grid cells to fill a large domain.
+To reduce blockage effects and computational cost, domain augmentation method use a large coarse grid to provide artificial boundary conditions to a small well-resolved domain [2].
 The outer flow field can also be augmented with a vortex particle-mesh method, reducing dispersive errors [3].
 Such methods enable tracking the wake for an extended time while enabling accurate modeling of the developing boundary layer on the object’s surface and can also include immersed boundaries [4].
-However, several (user-defined) parameters are required to correctly interpolate between the two domains, and the complexity of accurately coupling the domains makes theis apporach difficult to apply to existing flow solvers.
-For flows with one [6] or two unbounded directions [7, 8], coordinate transforms and/or analytical potential flow solution [9] can be applied in the unbounded directions.
-Potential flow boundary conditions are prescribed following the classical Helmholtz decomposition of the velocity field into an irrotational and an incompressible part
+However, these methods rely on a coupling of an external Lagrangian solver to an internal Eulerian solver formulated in different variables.
+An alternative approach is to utilize some form of potential flow boundary conditions, following the classical Helmholtz decomposition of the velocity field $\vec{u}(\vec{x})$ into an irrotational and an incompressible part
 $\vec{u}=\vec{u}_{\omega}+\nabla\phi+\vec{U}_{\infty}=-\nabla^{-2}\left(\nabla\times\vec{\omega}\right)+\nabla\phi+\vec{U}_{\infty}$ (1)
-where $\vec{u}_{\omega}$, $\nabla\phi$, $\vec{U}_{\infty}$ are the incompressible component, irrotational component, and the free-stream velocity, respectively.
-While these methods are elegant, they rely either on a single unbounded direction (as in [6] or formulating the problem in a coordinate system where two unbounded directions are transformed to a single one (for example, from Cartesian to Cylindrical in [8]), rendering them difficult to generalize to 3D flow problems.
-Other classes of methods solve unbounded Poisson equations by incorporating the unbounded conditions as Dirichlet boundary data, either by successive relaxed iterations of the Poisson problem with the Dirichlet data computed using the previous solution [10] or by using the James-Lackner [11, 12] algorithm that solves two nested domain Poisson problems, the inner-most with homogeneous Neumann boundary conditions and second outermost domain with inhomogeneous Neumann boundary conditions computed from a boundary convolution of the inner solution with the fundamental solution to the Poisson operator.
-To accelerate the convolution, multilevel approximation [13] or using local correction methods [14] can be used.
-Although these methods enforce unbounded boundary conditions, they require iterating a Poisson problem or solving it multiple times.
-Finally, fundamental solutions to discrete operators (Lattice Green’s functions or LGFs) and their asymptotic expansions can be used to solve the discrete Poisson equations in unbounded domains [15, 16].
-Points close to the source are evaluated with the discrete kernel (accurate), while points far away are evaluated with the asymptotic expansion (fast).
+where $\vec{u}_{\omega}$, $\nabla\phi$, $\vec{U}_{\infty}$ are the incompressible, irrotational, and the free-stream components, respectively.
+While these methods are elegant, they rely either on a single unbounded direction as in [5] or formulating the problem in a coordinate system where two unbounded directions are transformed to a single one (for example, from Cartesian to Cylindrical in [6]), which do not occur regularly in practical flow applications.
+A related class of methods solves unbounded Poisson equations by incorporating the unbounded conditions as Dirichlet boundary data, either by successive relaxed iterations of the Poisson problem with the Dirichlet data computed using the previous solution [7] or by using the James-Lackner [8, 9] algorithm that solves two nested domain Poisson problems, the inner domain with homogeneous Neumann boundary conditions and outer domain with inhomogeneous Neumann boundary conditions computed from a boundary convolution of the inner solution with the fundamental solution to the Poisson operator.
+To accelerate the convolution, multilevel approximation [10] or local correction methods [11] can be used.
+Finally, fundamental solutions to discrete operators (Lattice Green’s functions or LGFs) and their asymptotic expansions can be used to solve the discrete Poisson equations on unbounded domains [12, 13].
 These methods rely on the unbounded properties of the LGFs and can simulate unbounded viscous incompressible flow [14] and flows with immersed-boundaries [15].
-This allows for computational domains snug to the vorticity field, but adapting these methods to to existing Eulerian or immersed boundary solvers is difficult.
-In this manuscript, we propose a novel method for imposing external potential-flow boundary conditions of a viscous flow simulation by coupling to a vorticity-based description of the flow on the boundary via the Biot-Savart equation.
-The method retains the primitive variables formulation ($\vec{u}$ and $p$) of the governing equations, is compatible with immersed boundary methods and fast pressure projection methods, and does not require additional meshes to describe the flow.
+This allows for snug computational domains limited to regions of non-negligible vorticity or possibly a set of disjoint domains limited to active vorticity regions.
+Combined with a multi-resolution approach, the LGF approach can yield fast and accurate computation of 3D incompressible flows at a cost reduced compared to Cartesian meshes [16].
+However, the method relies on a structured grid and a constant coefficient Poisson equation, and the possibility of extending the method to unstructured mesh and variable coefficient Poisson equation problems (such as the one arising in [17]) is unclear.
+In this manuscript, we propose a novel method for imposing unbounded external flow boundary conditions on a finite Eulerian domain via a Biot-Savart integral solution applied on the domain’s boundary.
+The method retains the primitive variables formulation ($\vec{u}$ and the pressure $p$) of the governing equations, is compatible with immersed boundary methods, variable coefficient Poisson equations, fast pressure projection methods, and does not require additional meshes to describe the flow.
 In the next section, we describe the Biot-Savart boundary conditions and how they are injected into a projection algorithm to solve the Navier-Stokes equations.
-Next, we demonstrate the accuracy of the velocity reconstruction and the balance between speedup and accuracy in the reconstructed field.
-With confined vorticity and vortex wake examples, we demonstrate that the new boundary conditions significantly outperform standard reflective boundary conditions on significantly smaller domains.
+Next, we demonstrate the insensitivity of this unique approach to errors in the velocity reconstruction on the boundary, and use that to develop a simplified multi-level clustering approach for an optimal $O(N)$ Fast Multi-level Method to evaluate the Biot-Savart integral.
+We validate the correctness of the coupled equations using accelerating flow examples where the external flow is truly irrotational and demonstrate that the new boundary conditions vastly outperform standard reflective boundary conditions on significantly smaller domains.
 We present an unexpected result: the new method continues to perform well when the fundamental assumption of potential flow external to the domain is violated, and we use the unstable deflected wakes generated by fast flapping foils to study the limits of the approach.
@@
 Figure 1: Schematic of the fluid domain $\Omega$ with an immersed body $\mathcal{B}$ and their common (wet) interface $\partial\mathcal{B}$.
-The exterior of the computational domain $\Omega$ is denoted as $\partial\Omega$ with $\hat{n}$ the unit normal vector.
-We start by describing the governing equations and their numerical solution.
-The flow is governed by the Navier-Stokes equations for an incompressible flow.
-We discretize these equations on a staggered finite-volume mesh.
-The resulting system of equations is solved within a computational domain $\Omega(N)$, where $N$ is the number of degrees of freedom of the problem.
-The outer domain boundary is denoted at $\partial\Omega(N^{\cal S})$, where $\cal S$ is a factor linked to the dimensions of the problem.
-In 2D we have ${\cal S}=1/2$, and in 3D ${\cal S}=2/3$.
-The primary source of vorticity in the domain is an immersed body $\mathcal{B}$ with outer (wet) surface $\partial\mathcal{B}$ and unit normal vector $\hat{n}$, see Figure 1.
+The computational domain $\Omega$ has external boundary $\partial\Omega$ with $\hat{n}$ the unit normal vector.
+The flow is governed by the incompressible Navier-Stokes equations
+$\begin{array}[]{c}\rho\frac{\partial\vec{u}}{\partial t}+\rho\left(\vec{u}\cdot\nabla\right)\vec{u}=-\nabla p+\mu\nabla^{2}\vec{u}\\ \nabla\cdot\vec{u}=0\end{array}$ $\quad\quad\forall\ \vec{x}\in\Omega,$ (4)
+where $\rho,\mu$ are the fluid’s density and viscosity and $\Omega$ is the fluid domain , see Fig.
+1.
+These equations are supplemented with the no-slip boundary condition on the immersed body $\cal B$ with boundary $\partial\cal{B}$
+$\vec{u}=\vec{U}_{b}\quad\forall x\in\partial\cal{B}$ (5)
+and a velocity condition on the domain boundary $\partial\Omega$
+$\vec{u}(\vec{x})=f(\vec{x},\vec{U}_{\infty})$ $\quad\forall\ \vec{x}\in\partial\Omega,$ (6)
+where $\vec{U}_{b}$ is the body velocity and $f$ is the imposed condition and $\vec{U}_{\infty}$ is the far-field velocity.
+This work focuses on simulating external flow, i.e.
+$f$ is the identity function applied on an infinite domain $\partial\Omega\rightarrow\infty$.
+There are many possible conditions to apply on a finite domain to simulate the external conditions, and the simplest of these is a (local) reflection condition
+$\frac{\partial\vec{u}_{s}}{\partial\hat{n}}=0,\ \vec{u}_{n}=\vec{U}_{\infty}\cdot\hat{n}$ $\quad\forall\ \vec{x}\in\partial\Omega,$ (7)
+where we have decomposed the velocity field on the boundary in a boundary normal $\vec{u}_{n}$ and tangential part $\vec{u}_{s}$.
+We refer the reader to [1] for a discussion of pressure boundary conditions on the domain’s exterior.
+As mentioned previously, the boundary $\partial\Omega$ must be placed sufficiently far away from the body not to influence the decaying potential flow part of the velocity field and ensure enough space to properly advect the vorticity downstream.
+We propose to replace the typical “local” domain boundary conditions such as Eq.
+7 by a Biot-Savart integral over the vorticity inside the domain.
+A Green’s function solution to the vorticity equation Eq.
+1 gives the velocity at a point $\vec{x}$ as
+${\vec{u}}({\vec{x}})=\vec{U}_{\infty}+f(\vec{x};\vec{\omega},\Omega)=\vec{U}_{\infty}+\int_{\Omega}K_{n}({\vec{x}}-\vec{y})\times\vec{\omega}({\vec{y}})\text{ d}\vec{y}\quad\quad\forall\vec{x}\,\in\partial\Omega$ (8)
+where $K_{n}$ is the $n$-dimensional Biot-Savart kernel. This kernel takes the following form [18]
+$2D:\quad K(\vec{r})\equiv-\frac{\vec{r}}{2\pi|\vec{r}|^{2}},\qquad\qquad 3D:\quad K(\vec{r})\equiv-\frac{\vec{r}}{4\pi|\vec{r}|^{3}}.$ (9)
+The vortex-induced velocity of the Biot-Savart integral is simply added to the uniform background velocity $\vec{U}_{\infty}$ to determine the total velocity at any point.
+We will omit further discussion of the trivial uniform flow contribution in this section.
+This formulation implicitly assumes the flow external to the domain is irrotational and that the flow is incompressible everywhere.
+This assumption is met when all the vorticity is confined within the computational domain such as for the flow around accelerated bodies when the time $\frac{tU}{L}\lesssim t^{\text{crit.}}$, with $t^{\text{crit.}}$ the critical time for vortex detachment [19].
+However, we will show that the results of using Eq.
+8 in an Eulerian solver are still highly accurate for flows with wakes that extend to the domain boundary.
+Note that on a staggered grid, we only need to apply the integral equation (Eq.
+8) to compute the velocity normal to the domain face.
+For ghost cells external to the computational domain, we use the local derivative conditions $\nabla\cdot\vec{u}=\nabla\times\vec{u}=0$ which are consistent with the assumption that the flow is potential outside the domain.
+Specifically, once we have used Eq.
+8 to set $u_{n}$ along the domain boundary we apply the zero curl condition to all tangential velocity components in the ghost cells and then the zero divergence condition to fill the remaining normal component each ghost cell.
+For example, in 2D we sequentially apply
+$\frac{\partial\vec{u}_{s}}{\partial\hat{n}}=\frac{\partial\vec{u}_{n}}{\partial\hat{s}}\quad\text{and}\quad\frac{\partial\vec{u}_{n}}{\partial\hat{n}}=-\frac{\partial\vec{u}_{s}}{\partial\hat{s}}.$
+This process can continue for multiple layers of ghost cells, marching away from the domain boundary.
+There are two primary issues with the use of Eq.
+8 as the domain boundary condition for the incompressible Eulerian solver: (i) projecting the velocity into a divergence-free field consistent with the Biot-Savart equation, and (ii) the fast evaluation of Eq.
+8.
+We demonstrate in the following section that these issues may be efficiently and accurately overcome largely without modifying the method used to solve the Navier-Stokes equation.
+2.1 Biot-Savart Projection Algorithm
 We use the standard projection scheme [20] to solve the coupled velocity-pressure system inside a predictor-correct method to achieve second-order temporal convergence of the flow variables [17].
-To simplify the discussion, we present the method for the first step of the predictor-corrector; generalization to the second-order Heun’s corrector method or higher-order time integration method is straightforward.
+To simplify the discussion, we present the method for the first step of the predictor-corrector; generalization to the second-order Heun’s corrector or higher-order time integration method is straightforward.
 The projection method starts with an explicit estimate of the intermediate velocity field $u^{*}$
@@
 Here we will focus on fully explicit methods.
-On the immersed body $\cal B$, the intermediate velocity is subject to the condition
-$\vec{u}^{*}=\vec{U}_{b}\quad\forall x\in\partial\cal{B}.$ (3)
-where $\vec{U}_{b}$ is the body velocity.
-Body-fitted methods enforce this condition directly, while immersed boundary methods enforce it indirectly; either by reformulating the governing equations near the boundary such as [21, 22] or by modifying the discretization near $\cal{B}$ [23].
-Additionally, Eq. (2) must be supplemented with far-field boundary conditions on the domain’s outer boundaries. The simplest of these is a reflection condition
-$\frac{\partial\vec{u}^{*}_{s}}{\partial\hat{n}}=0,\ \vec{u}^{*}_{n}=U_{n}$ $\quad\forall\ \vec{x}\in\partial\Omega(N^{\cal S}),$ (4)
-where we have decomposed the velocity field on the boundary in a boundary normal $\vec{u}^{*}_{n}$ and tangential part $\vec{u}^{*}_{s}$.
-We refer the reader to [1] for a discussion of pressure boundary conditions on the domain’s exterior.
 Projection of the divergent velocity field onto the solenoidal space is achieved by computing a pressure field $p$ that removes the divergent part of the intermediate velocity field
@@
 $\frac{\partial\vec{u}^{t+\Delta t}_{s}}{\partial\hat{n}}=0,\ \vec{u}^{t+\Delta t}_{n}=U_{n}$ $\quad\forall\ \vec{x}\in\partial\Omega(N^{\cal S}).$ (13)
-As mentioned previously, the boundary $\partial\Omega$ must be placed sufficiently far away from the body not to influence the fast decaying (potential flow) part of the velocity field and ensure enough space to properly advect the vorticity downstream.
-2.1 Biot-Savart boundary conditions
-We propose to replace the typical “local” domain boundary conditions such as Eq.
-(4) &  (7) by a Biot-Savart integral over the vorticity inside the domain.
-A Green’s function solution to the vorticity equation Eq.
-(1) gives the velocity at a point $\vec{x}$ as
-${\vec{u}}({\vec{x}})=f(\vec{x};\vec{\omega},\Omega)=\int_{\Omega}K_{n}({\vec{x}}-\vec{y})\times\vec{\omega}({\vec{y}})\text{ d}\vec{y}\quad\quad\forall\vec{x}\,\in\partial\Omega(N^{\cal S})$ (8)
-where $K_{n}$ is the $n$-dimensional Biot-Savart kernel. This kernel takes the following form [24]
-$2D:\quad K(\vec{r})\equiv-\frac{\vec{r}}{2\pi|\vec{r}|^{2}},\qquad\qquad 3D:\quad K(\vec{r})\equiv-\frac{\vec{r}}{4\pi|\vec{r}|^{3}}.$ (9)
-This formulation implicitly assumes the flow external to the domain is irrotational and that the flow is incompressible everywhere.
-This assumption is met when all the vorticity is confined within the computational domain such as for the flow around accelerated bodies when the time $\frac{tU}{L}\lesssim t^{\text{crit.}}$, with $t^{\text{crit.}}$ the critical time for vortex detachment [25].
-However, we will show that the results of using equation (8) in an Eulerian solver are still highly accurate for flows with wakes that extend to the domain boundary.
-Note that on a staggered grid, we only need to apply the integral equation (8) to compute the velocity normal to the domain face.
-For ghost cells external to the computational domain, we use the local derivative conditions $\nabla\cdot\vec{u}=\nabla\times\vec{u}=0$ to compute the velocity.
-These conditions are consistent with the assumption that the flow is potential outside the domain and fast to apply.
-Specifically, we apply the zero curl condition to all tangential velocity components in the ghost cells and then the zero divergence condition to the normal components in the ghost cells, marching away from the domain boundary.
+The main contribution of this manuscript is injecting the Biot-Savart integral into the classical projection scheme and solving the resulting coupled problem.
+The momentum update using the new Biot-Savart domain condition proceeds as before: forming the intermediate velocity with Eq.
+10 and projecting it to be divergence free with Eq.
+12.
+However, the new boundary condition introduces a coupling between the body and domain boundaries.
+To see this, we substitute the update equation for the velocity at the new time-step (Eq.
+12) into the Biot-Savart equation for the boundary velocity (Eq.
+8) to give to complete equation for the boundary velocity at $t+\Delta t$
+$\vec{u}^{t+\Delta t}(\vec{x}_{b})=f(\vec{x_{b}},\nabla\times\vec{u}^{t+\Delta t},\Omega)=f(\vec{x_{b}},\nabla\times\vec{u}^{*},\Omega)-f\left(\vec{x_{b}},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega\right)$ (14)
+While the gradient of a field cannot induce curl in an unbounded fluid region (×$\nabla\times\nabla\phi\equiv 0$), the pressure generates a thin layer of vorticity on the body boundary $\partial\cal B$ every time-step [21].
+This means that the last term in Eq.
+14 is non-zero, introducing a non-local coupling between every point on the body boundary and every point on the domain boundary.
+Because of the pressure contribution in Eq. 14, the pressure projection (Eq. 11) becomes
+$\nabla\cdot\left[\frac{\Delta t}{\rho}\nabla p+f\left(\vec{x_{b}},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega\right)\right]=\nabla\cdot\left[\vec{u}^{*}+f(\vec{x_{b}},\nabla\times\vec{u}^{*},\Omega)\right],\quad\forall\ \begin{array}[]{l}\vec{x}\in\Omega(N),\\ \vec{x}_{b}\in\partial\Omega(N^{\mathcal{S}})\end{array}$ (15)
+The linear operator resulting from discretizing Eq.
+11 with a local finite-difference finite-volume or finite-element method has only $O(N)$ non-zero terms, but the coupling induced by the Biot-Savart equation results in a dense matrix with at least $O(N^{1+S})$ non-zero terms.
+Note that this coupling is similar to the additional terms present in the Poisson equation when correctly imposing pressure boundary conditions for immersed boundary methods; see, for example, [22, 17].
+Failing to correctly enforce the boundary conditions during projection leads to large errors in the predicted pressure and forces.
+We must accurately include the coupling term in Eq.
+14, but we want to avoid constructing the dense matrix or requiring substantial changes to existing Poisson solvers to make it practical for Eulerian schemes to use Biot-Savart boundary conditions.
+This is possible through a iterative operator-split method akin to the James-Lackner algorithm for potential flows
+$\nabla\cdot\frac{\Delta t}{\rho}\nabla p^{k+1}=\nabla\cdot\left[\vec{u}^{*}+f\left(\vec{x_{b}},\nabla\times\left[\vec{u}^{*}-\frac{\Delta t}{\rho}\nabla p^{k}\right],\Omega\right)\right],\quad\forall\ \begin{array}[]{l}\vec{x}\in\Omega(N),\\ \vec{x}_{b}\in\partial\Omega(N^{\mathcal{S}}),\end{array}$ (16)
+using the previous pressure solution $p^{k}$ to treat the Biot-Savart term as a residual and iterating until both the incompressible and boundary conditions are satisfied.
+Eulerian solvers typically use iterative methods to determine the pressure field in the projection step, making this a natural way to incorporate the coupling between the domain and body boundaries, without modifying the Poisson solver itself.
+In this work, we use a multigrid solver for the Poisson system, as described in [23], and updating the domain velocities and corresponding Poisson residuals at the end of each V-cycle is sufficient to ensure that the final pressure and velocity field without increasing the number of iterations.
+Our integration of Eq.
+16 into the unsteady momentum update significantly improves the stability of the coupled system compared to similar applications in potential flow such as  [7].
+The error introduced by the operator partitioning is bounded by the change in pressure on the body over a single time step.
+We therefore never require relaxation to achieve convergence, even when impulsively starting a flow from rest with domain boundaries a small fraction of the body length away from the body boundaries.
+In addition, the resulting coupled system is remarkably robust to errors in the velocity reconstruction because the Biot-Savart equation is only used to set the boundary conditions, not update the entire flow field as in vortex-based methods.
+Consider the effect of neglecting the contribution from a vorticity source $\Delta\omega$ that has just left the domain, meaning this contribution is missing from ×$\nabla\times\vec{u}^{*}$ in Eqs.14 and 16.
+As only the boundary velocities are updated in Eq.
+14, the momentum errors in the local Eulerian Navier-Stokes update are limited to the downwind characteristics from the domain boundary, and the impact on the (critical) inflow boundary regions are typically the smallest, as they are the farthest from the vorticity leaving the outflow boundary regions.
 2.2 Fast evaluation of the Biot-Savart integral
-$\mathcal{D}^{(4)}$
+The aim of Biot-Savart boundary conditions is to reduce the computational cost compared to a large grid by providing an accurate estimate of external flows using small domains.
+While the boundary conditions are agnostic of the method used to compute the integral in Eq.
+8, a naive convolution over $N$ vorticity “source” locations $\vec{y}$ would require $O(N)$ operations for every “target” velocity $\vec{u}(\vec{x})$ on the boundary.
+The overall $O(N^{1+\cal S})$ cost would likely overwhelm any benefit of the reduced grid size, motivating a faster Biot-Savart evaluation approach.
+Highly precise Fast Multipole Methods (FMMs) have been developed to a reduce this cost to $O(\log N)$ or even $O(1)$ operations per target, [24, 25, 26].
+Our context allows for significant specialization and simplification to a standard FMM.
+(i) A general FMM considers a general set of source locations, and they are typically optimized to evaluate on the same set of targets.
+In contrast, our sources are equally spaced over the domain and our targets are equally spaced on it’s boundary, meaning we have only $O(N^{\cal S})$ targets which are exclusively exterior to the source set.
+(ii) The high error tolerance discussed in the section above means truncating the multipole has no effect on the satisfaction of the divergence-free constraint (set by the Poisson solver tolerance) and little effect on the solution accuracy overall.
 $\mathcal{D}^{(4)}$
@@
 $\Omega$
-$\mathcal{B}$
 $\vec{y}\,^{(2)}$
@@
 ${S}\,^{(2)}$
-Figure 2: Schematic of the multilevel approach to evaluate the Biot-Savart integral for a boundary point $\vec{x}$ over the nested domains $\mathcal{D}^{(1)}\cup\mathcal{D}^{(2)}\cup\mathcal{D}^{(3)}\cup\mathcal{D}^{(4)}$.
+$\mathcal{P}^{(3)\to(4)}$
+$\mathcal{D}^{(2)}$
+$\mathcal{B}$
+Figure 2: Schematic of the multilevel approach to evaluate the Biot-Savart integral for a boundary point $\vec{x}$ over the nested subdomains $\mathcal{D}^{(1)}\cup\mathcal{D}^{(2)}\cup\mathcal{D}^{(3)}\cup\mathcal{D}^{(4)}$.
 The second domain half-width is shown, ${S}^{(2)}$.
 The computational domain $\Omega$ and its boundary $\partial\Omega$ and the immersed body $\mathcal{B}$ are shown.
-We also show a schematic of the coarsening of the grid between the domains.
-Domain and cell size are only for representation proposes.
-The aim of this novel boundary conditions is to reduce the computational cost compared to a large grid by providing an accrate estimate of the far field conditions.
-A naive evaluation of Eq.
-8 would require $O(N)$ operations for each boundary point, meaning $O(N^{1+\cal S})$ cost overall, potentially overwhelming any savings from reducing the size of the domain.
-In vortex particle-mesh methods, the integral is usually evaluated with Fast Multipole Method that allows reducing this cost to $O(\log N)$ operations.
-Here, we use a multilevel algorithm to achieve the same speed-up in our Eulerian setting.
-We construct the multilevel Biot-Savart operator for each point $\vec{x}$ on the domain boundary $\partial\Omega$ by representing the complete domain $\Omega$ with a combination of gradually coarser meshes
+We also show a schematic of the coarsening of the grid between the domains and the multilevel pooling $\mathcal{P}^{(3)\to(4)}$ of the cell’s circulation (circular-colored markers).
+Based on this context, we use a simple multi-level vorticity source clustering method which leverages their uniform distribution in our Eulerian domain.
+We construct a multi-level vorticity field $\omega^{(i)}$ where level $i=1$ is the finest and $l$ is the coarsest and the cell size doubles with each level $h^{(i+1)}=2h^{(i)}$.
+The levels are filled from finest to coarsest recursively by uniformly pooling the circulation $\Gamma=\omega h^{2}$ on the level above
+$\Gamma^{(i+1)}=\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}$ (17)
+where the pooling $\mathcal{P}$ for a cell at level $i+1$ simply sums over the corresponding sub-cells in level $i$, see Fig.
+2.
+Thus, our multi-level field conserves the total circulation to machine precision, as recommended by [2] and others.
+Filling the entire multi-level source field requires $O(N)$ operations.
+In this work, we present results from two boundary velocity reconstruction approaches using the multi-level source clustering described above:
+  1. 1.
+    tree-sum: In the first approach, the integral over the original domain $\Omega$ in Eq. 8 is replaced with a sum over the contribution from a sequence of subdomains
     $\Omega=\mathcal{D}^{(1)}(\vec{x})\cup\mathcal{D}^{(2)}(\vec{x})\cup\cdots\cup\mathcal{D}^{(l-1)}(\vec{x})\cup\mathcal{D}^{(l)}(\vec{x}),$ (18)
-where level $1$ is the finest domain and $l$ is the coarsest, and each domain is non-overlapping and their union fully covers $\Omega$, Fig 2.
-In this work, we use nested rectangular boxes of half-width $S^{(i)}$ centered on $\vec{x}$, and scale this width with the cell size which doubles at each level, i.e.
-$h^{(i)}=h^{(1)}2^{i-1}$.
-Therefore the domains are determined by $\vec{x}$ and a non-dimensional half-width
-$\tilde{S}=\frac{{S}^{(l)}}{2^{l-1}},$ (11)
-which is equal on every level.
-This domain decomposition is chosen so that the contributions to Eq.
-(8) are more accurately calculated when close to $\vec{x}$, and the smaller contributions further from $\vec{x}$ are quickly estimated on the coarser levels.
-With this set of $l$ non-overlapping domains $\mathcal{D}^{(i)}$, the integral in Eq. 8 can be replaced with a sum over all the contribution from the domains
-$f(\vec{x};\vec{\omega},\Omega)\approx\sum_{i=1}^{l}f(\vec{x};\vec{\omega}^{(i)},\mathcal{D}^{(i)}).$ (12)
-Once the multilevel field $\omega^{(i)}$ is computed, the total number of operations to evaluate Eq.
-12 for a simulation of dimension $d=2$ or $3$ is therefore $O(l\tilde{S}^{d})=O(\log N)$ since $\tilde{S}$ is constant and $l=O(\frac{1}{d}\log N)$.
-In practice, we use the circulation over a cell face $\Gamma=\omega h^{2}$ instead of the vorticity in the Biot-Savart formulation.
-The circulation on a coarser level is calculated by uniformly pooling the circulation on the level above
-$\Gamma^{(i+1)}=\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}$ (13)
-where $\Gamma^{(1)}$ is the original finest level field and the pooling $\mathcal{P}$ for a cell at level $i$ simply sums over the corresponding subcells in level $i-1$, see [26].
-Thus, our multi-level field $\Gamma^{(i)}$ conserves the total circulation to machine precision, as recommended by [2] and others.
-Finally, we derive the error induced by this multilevel approach, starting with the 2D-case and extending the result to 3D.
-The velocity at location $\vec{x}$ induced by the circulation from a single cell on level $i+1$ is given by (see Fig.
-2)
-$\vec{u}^{(i+1)}(\vec{x})=\frac{\vec{r}}{2\pi|\vec{r}|^{2}}\times\vec{e}_{z}\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}$ (14)
-where $\vec{r}=\vec{y}-\vec{x}$, $y$ is the center of the cell, and $\vec{e}_{z}$ is the out-of-plane unit vector. The velocity induced by the four corresponding cells on level $i$ is
-$\vec{u}^{(i)}(\vec{x})=\mathcal{P}^{(i)\to(i+1)}\left(\frac{\vec{r}_{j}}{2\pi|\vec{r}_{j}|^{2}}\times\vec{e}_{z}\Gamma^{(i)}_{j}\right),$ (15)
-where $\vec{r}_{j}=\vec{y}_{j}-\vec{x}$ and $y_{j}$ are the four cell centers.
-We can further decompose this expression by noting that $\vec{r}_{j}=\vec{y}-\vec{x}-\vec{\delta r}_{j}=\vec{r}-\vec{\delta r}_{j}$.
-For sufficiently large $\vec{r}$ we can assume that $|\vec{r}|\approx|\vec{r}-\vec{\delta r}_{j}|$ which allows us to separate the kernel’s influence into two parts
-$\vec{u}^{(i)}(\vec{x})=\mathcal{P}^{(i)\to(i+1)}\left(\frac{\vec{r}}{2\pi|\vec{r}|^{2}}\times\vec{e}_{z}\Gamma^{(i)}_{j}-\frac{\vec{\delta r}_{j}}{2\pi|\vec{r}|^{2}}\times\vec{e}_{z}\Gamma^{(i)}_{j}\right).$ (16)
-After rearrangement, the first term equals Eq. (14). The second term is the error due to the multilevel approximation. Its magnitude is
-$\varepsilon(\vec{x})^{(i)}=\mathcal{P}^{(i)\to(i+1)}\frac{|\vec{\delta r}_{j}|}{2\pi|\vec{r}|^{2}}\Gamma^{(i)}_{j},$ (17)
-as $\vec{e}_{z}$ is a unit vector.
-On uniform Cartesian meshes, $|\vec{\delta r}_{j}^{(i)}|\sim 2^{l}$, while $|\vec{r}|^{2}\sim(2^{l}\tilde{S})^{2}$.
-An upper bound for the total error in the induced velocity at a point $\vec{x}$ due to the multilevel evaluation of the Biot-Savart integral is therefore
-$\varepsilon(\vec{x})\approx\frac{1}{2\pi\tilde{S}^{2}}\sum_{i=2}^{l}\frac{1}{2^{i}}\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}.$ (18)
-which decays as $\tilde{S}^{-2}$.
-The content of each sum is a decreasing fraction of the total vorticity contained in the domain $\mathcal{D}^{(i)}$.
-As expected, taking $\tilde{S}>\text{Domain}$ yields $\varepsilon(\vec{x})=0$, and as $\Gamma^{(i)}\to 0,\varepsilon(\vec{x})\to 0$, finally as $\tilde{S}\to 0,\varepsilon(\vec{x})\to\infty$.
-Similarly, in 3D the upper bound for the error is given by
-$\varepsilon(\vec{x})\approx\frac{1}{4\pi\tilde{S}^{3}}\sum_{i=2}^{l}\frac{1}{2^{2i}}\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}.$ (19)
-which decays as $\tilde{S}^{-3}$.
-Similarly, the content of the sum is a (even smaller) decreasing fraction of total vorticity contained in each level, improving the method’s accuracy in 3D compared to 2D.
-2.3 Biot-Savart & Projection Algorithm
-The momentum update using the new Biot-Savart domain condition proceeds as before: forming the intermediate velocity with Eq.
-2 and projecting it to be divergence free with Eq.
-6.
-However, the new boundary condition introduces a coupling between the body and domain boundaries.
-To see this, we write the equation for the boundary velocity on $\vec{x}_{b}\in\partial\Omega(N^{\cal S})$ at the new time-step as
-$\vec{u}^{t+\Delta t}(\vec{x}_{b})=f(\vec{x_{b}},\nabla\times\vec{u}^{t+\Delta t},\Omega)=f(\vec{x_{b}},\nabla\times\vec{u}^{*},\Omega)-f\left(\vec{x_{b}},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega\right)$ (20)
-where we’ve substituted Eq.
-6.
-The last term in Eq.
-20 is non-zero on cells touching the body boundary $\partial\cal B$ because, while the gradient of a field cannot induce curl in an unbounded fluid region (×$\nabla\times\nabla\phi\equiv 0$), the pressure boundary condition on the immersed body generates a thin layer of vorticity on $\partial\cal B$.
-This contribution is unknown until after the Poisson system of Eq.
-5 is solved, but updating the normal velocity on $\partial\Omega(N^{\cal S})$ will change the source term for Eq.
-5 on cells touching the domain boundary, requiring a new pressure solution.
-This coupling between the domain and body boundaries results in a mildly non-linear fixed-point system.
-This weak non-linearity can be solved efficiently with a deferred correction approach, which explicitly updates the source term of the Poisson problem during pressure solver iteration.
-We use a Multi-Grid solver for the Poisson system, as described in [26], and updating the domain velocities and corresponding Poisson residuals at the end of each V-cycle was sufficient to ensure the final pressure and velocity field satisfy both the momentum and divergence-free conditions without increasing the number of solver iterations.
-Note that this coupling is only apparent in immersed boundary methods if the pressure boundary conditions are correctly imposed on the immersed body during the projection step; see for example [27, 20].
-This approach differs from the one in [10] for unbounded Poisson problems with embedded Dirichlet boundaries as we directly adjust the residual of the (iterative) Poisson solver and we do not iterate over the coupled Poisson/interior boundary problem.
-It is also different to the James-Lackner algorithm for the same reason [13].
-Lastly, we note that the small multilevel error discussed in the previous section means that global mass conservation is no longer guaranteed to machine precision on the domain boundary.
-As such, we apply a uniform correction to the velocity computed with Eq.
-20
-$\vec{u}(\vec{x}_{b})=\vec{u}(\vec{x}_{b})-\frac{\oint_{\partial\Omega}\hat{n}\cdot\vec{u}\text{ d}S}{\oint_{\partial\Omega}\text{ d}S}.$ (21)
-ensuring a globally divergence-free field is passed to the pressure projection method on each iteration.
-3 Multilevel reconstruction validation
-We validate the multilevel reconstruction algorithm presented above using 2D and 3D analytic vorticity fields.
-We also study the multilevel reconstruction error and the speedup generated by selecting the correct kernel radius from the above analysis.
-Our Biot-Savart boundary conditions are implemented inside WaterLily.jl [^2] https://github.com/weymouth/WaterLily.jl, a fast, immersed-boundary flow solver based on the Boundary data immersion method [22] that can execute both on CPU and GPU [28].
+    where we define the sub-domains as nested rectangular boxes of half-width $S^{(i)}=h^{(i)}\tilde{S}$ centered on the target location $\vec{x}$, Fig 2.
+As the sub-domains are non-overlapping and fully cover $\Omega$, Eq.
+8 becomes
+    $\vec{u}(\vec{x})\approx\sum_{i=1}^{l}f(\vec{x};\vec{\omega}^{(i)},\mathcal{D}^{(i)}(\vec{x})).$ (19)
+    This corresponds to a direct sum for each target $\vec{x}$ over an oct-tree source field in 3D (or a quad-tree field in 2D).
+Therefore, we call this the “tree-sum” approach.
+As each subdomain has $O(\tilde{S}^{n})$ sources and the number of levels $l=O(\log N)$, the tree-sum requires $O(\tilde{S}^{n}\log N)$ operations per target.
+  2. 2.
+    $\ell$M: In the second approach, we also cluster the domain boundary target points $\vec{x}^{\,(i)}$, doubling the spacing at each level.
+We then define the level-$i$ kernel interaction velocities as
+    $\vec{v}^{\,(i)}(\vec{x}^{\,(i)})=f(\vec{x}^{\,(i)};\vec{\omega}^{(i)},\mathcal{D}^{(i)}(\vec{x}^{\,(i)})).$ (20)
+    The target velocities are then unpooled recursively from the coarsest level to the finest
+    $\vec{u}^{\,(i)}\approx\vec{v}^{\,(i)}+\mathcal{P}^{(i+1)\to(i)}\vec{u}^{\,(i+1)}$ (21)
+    where $u^{\,(l+1)}=0$ and the unpooling operator is bi-linear interpolation to the target locations $\vec{x}^{\,(i)}$.
+As the interaction velocities are independent at each level, they require a constant $O(\tilde{S}^{n})$ operations per target.
+Therefore, we call this the Fast Multi-level Method ($\ell$M).
+A key difference between these approaches and a classic FMM is that the contributions from each clustered source point are obtained by evaluating Eq.
+9 directly and not through a classical or kernel-independent multipole expansion, as done in [25, 12].
+As such, the domain half-width $\tilde{S}$ controls the reconstruction error and A proves that this error is bounded and inversely proportional to the number of subdomain points $\tilde{S}^{n}$.
+The results sections below demonstrate that these approaches are more than sufficient for high-accuracy external flow predictions.
+Finally, we note that we can parallelize the velocity reconstruction to further accelerate the method since each Biot-Savart kernel evaluation is independent.
+Our implementation applies GPU or CPU multi-threading to parallelize the calculation and pooling of the vorticity across the sources, as well as the kernel function evaluation and velocity unpooling across the targets.
+3 Confined vorticity applications
+We first validate the application of the Biot-Savart boundary conditions using a set of impulsive flow problems.
+The flow is truly irrotational outside the computational domain, matching the implicit assumptions required to use the Biot-Savart equations to reconstruct the velocity on the boundary.
+Our Biot-Savart boundary conditions are implemented inside WaterLily.jl [^1] https://github.com/WaterLily-jl/WaterLily.jl, a fast, immersed-boundary flow solver based on the Boundary data immersion method [27] that can execute both on CPU and GPU [28].
 To leverage the GPU capability of the flow solver and fully benefit from our novel boundary conditions, we perform all computations presented herein on an NVIDIA-A6000 GPU (48GB memory).
 The complete code and instructions needed to reproduce these results are freely available at https://github.com/weymouth/BiotSavartBCs.jl.
-3.1 2D validation: Lamb Chaplygin vortex-dipole
-The velocity field generated by the Lamb-Chaplygin dipole is defined by the scalar stream function
-$\psi(r)=\begin{cases}\frac{-2UJ_{1}(\beta r)}{\beta J_{0}(\beta R)},\quad\text{for}\quad r<R,\\ U\left(\frac{R^{2}}{r}-r\right),\quad\text{for}\quad r\geq R,\end{cases}$ (22)
-where $J_{0}$ and $J_{1}$ are the zeroth and first Bessel functions of the first kind, respectively, and $\beta$ is a parameter with the value $\beta=1.2197\pi/R$.
-The 2D velocity field is given by $\vec{u}=-\vec{e}_{z}\times\nabla\psi$, and the core of the vortex is defined by the isoline $\psi(r=R)=0$.
-$10^{-1}$
-$10^{-2}$
-$10^{-3}$
-$10^{-4}$
-$10^{-5}$
-$10^{-6}$
-$a)$ $S=2$
-$c)$ $S=8$
-$d)$ $S=16$
-$f)$ $S=128$
-$b)$ $S=4$
-$e)$ $S=32$
-Figure 3: $L_{2}$-norm of the error in reconstructed velocity field outside the core of the 2D Lamb-Chaplygin vortex for varying (multi-level) kernel size.
-We reconstruct the velocity field outside the vortex core using the internal vorticity and our multilevel Biot-Savart method and compare that to the analytic result.
-We reconstruct the velocity on a domain of size $[2^{8},2^{8}]$.
-The dependency of the absolute error in the reconstructed velocity field on the radius of the multilevel refinement criterion are shown in Fig.
-3.
-Setting the refine size $S$ to the domain size $\cal D$ eliminates the multi-level-error, but the result is not exact since the analytic vorticity distribution is sampled on a finite grid of size $\Delta x=0.05R$.
-This error depends primarily on the distance from the vortex, Fig.
-3(f).
-Progressively reducing the refinement size $S$ introduces the multilevel error quantified in the previous section, which is only weakly dependent on the distance to the vortex.
-This is further quantified in Fig.
-4(a), which shows the maximum error in the reconstructed velocity field at a distance $d$ from the vortex from the Lamb-Chaplygin vortex.
-The size $S$ is shown to dominate the reconstruction error, with fairly uniform total error levels once $S\leq 2^{3}=8$.
-Fig.
-4(b) reports the measured computational speed-up achieved by the multilevel method relative to setting $S=\Omega$, demonstrating up to 300$\times$ acceleration.
-Figure 4: Maximum error in the reconstructed velocity field generated by a 2D Lamb-Chaplygin vortex dipole for various kernel sizes and distances from the vortex core (left).
-Corresponding speed-up obtained by reducing the kernel size for the multilevel algorithm (right).
-The mesh has dimensions $[2^{8},2^{8}]$.
-3.2 3D validation: Hill vortex
-This validation is repeated in 3D using the classical Hill vortex, whose radial and tangential velocity components are given by
-$v_{r}$ $=\frac{1}{r^{2}\sin\theta}\frac{\partial\psi}{\partial\theta},\qquad v_{\theta}=-\frac{1}{r\sin\theta}\frac{\partial\psi}{\partial r}.$ (23)
-and the stream function
-$\psi(r,\theta)=\begin{cases}-\frac{3U}{4}\left(1-\frac{r^{2}}{R^{2}}\right)r^{2}sin^{2}\theta\quad\text{for}\,r\leq R,\\ \,\,\frac{U}{2}\left(1-\frac{r^{3}}{R^{3}}\right)r^{2}sin^{2}\theta\quad\text{for}\,r>R.\end{cases}$ (24)
-with the corresponding errors and speed-up shown in Figure 5.
-Note that the analytic vorticity of the hill vortex is discontinuous, increasing the truncation error near the vortex compared to the smooth 2D Lamb-dipole.
-As argued in Sec.
-2.3 the multilevel error decreases extremely quickly as $S$ is increased for three-dimensional flows.
-This is due to the scaling of the Biot-Savart kernel itself.
-The 3D case demonstrates a stronger sensitivity to the distance to vortex, as expected from the error estimate.
-Similarly, the speed-up is much faster in 3D because each doubling gathers 8 cells instead of 4, with a max observed speed-up of nearly 20,000$\times$ using the multilevel method.
-Figure 5: Maximum error in the reconstructed velocity field generated by a 3D Hill vortex for various kernel sizes and distances from the vortex core (left).
-Corresponding speed-up obtained by reducing the kernel size for the multilevel algorithm (right).
-The mesh has dimensions $[2^{8},2^{8},2^{8}]$.
-4 Confined vorticity applications
-Next, we validate the application of the Biot-Savart conditions to a set of impulsive flow problems.
-The flow is truly irrotational outside the computational domain, matching the implicit assumptions required to use the Biot-Savart equations to reconstruct the velocity on the boundary.
+3.1 Flow around a 2D cylinder at $\text{Re}=550$
+We start our analysis with the flow behind a 2D circular cylinder of diameter $D$.
+The cylinder is immersed in a viscous fluid and impulsively accelerated to free-stream velocity $U$.
+The viscosity is set such that the Reynolds number is $\text{Re}=\frac{UD}{\nu}=550$.
+We set $D=128$ cells for all tests and use a $2\times 1$ rectangular domain, varying the width $W$ from $=2D\ldots 24D$.
+Figure 3(left) shows the vorticity field on the smallest domain at convective time $t^{*}=tU/D=4$.
+Figure 3(right) compares the predicted drag coefficient using the new Biot-Savart boundary conditions to the analytical solution [29] (for $t^{*}\ll 1$) and two sets of reference data from high-resolution vortex-based methods [30, 31].
+The new condition is able to accurately predict the total drag force on the cylinder, matching the theoretical and vortex-based methods even when using the smallest $4D\times 2D$ domain with 50% blockage.
+Reflective boundary conditions induce extremely large blockage effects for small domains and only slowly converge to the external flow solution.
+Even using a domain 100 times larger produces noticeable deviations for $t^{*}>3.5$.
+Figure 3: Impulsively started circular cylinder at $\text{Re}=550$ validation case.
+(Right) Vorticity field $\omega D/U$ at convective time $tU/D=4$ over the full $2D\times 4D$ domain.
+(left) Early time history of the drag coefficient $C_{d}=F_{d}/\frac{1}{2}\rho U^{2}D$ for various domain blockage ratios $D/W$.
+The present Biot-Savart BCs match the theoretical and vortex-based methods even with 50% blockage.
 3.2 2D and 3D flow around an accelerating disk
 The flow around a 2D or 3D disk accelerated from rest in a quiescent flow is purely potential for $t\to 0^{+}$.
-As a result, an analytical expression for the added mass coefficient can be used to validate our novel potential flow boundary condition when injected in a moment step of the Navier-Stokes equations.
+As a result, an analytical expression for the added mass coefficient can be used to further validate our novel potential flow boundary condition when injected in a moment step of the Navier-Stokes equations.
 The disk of radius $R$ and diameter $D=2R$ is given a constant acceleration $a$ for the initial non-dimensional time $t^{*}=at/U\leq 1$ and is given a constant speed after this, ie
 $U(t)=\begin{cases}at,\quad\text{if }at/U\leq 1,\\ U,\quad\text{else}.\end{cases}$ (22)
-We perform 2D and 3D simulations of the flow around a plate/circular disk immersed in a domain of dimension $1.5D\times 1.5D(\times 1.5D)$.
+We perform 2D and 3D simulations of the flow around a plate/circular disk immersed in a domain of dimension $1.5D\times 1.5D\,\,(\times 1.5D)$ with a resolution of 86 cells per diameter ($D$).
 From potential flow theory, the added-mass coefficient for an impulsively started flow is
@@
 We note here that the 2D and 3D problems possess axes of symmetry that would allow only a portion of the domain to be modeled.
-This can be easily accounted for in the Biot-Savart boundary conditions by adding the contributions of the images of each vortex cell, see A.
+This can be easily accounted for in the Biot-Savart boundary conditions by adding the contributions of the images of each vortex cell, see B.
 However, to avoid unnecessary complications in this manuscript, we model the full plate and do not use the problem’s symmetries.
-Figure 6 shows the vorticity and pressure field at three consecutive times during the disk’s motion on the midplane of the 3D circular disk.
+Figure 4 shows the vorticity and pressure field for three consecutive times during the disk’s motion on the midplane of the 3D circular disk.
 The top row shows the results with the standard reflective boundary conditions (the entire computational domain is shown), and the bottom row shows the flow obtained with the Biot-Savart boundary conditions.
@@
 A slice through the middle of the entire computational domain is shown.
-The top half shows 10 iso-contour of the vorticity, equally spaced in the interval $\omega L/U\in\pm 20$.
+The top half shows 10 iso-contour of the vorticity, equally spaced in the interval $\omega R/U\in\pm 20$.
 The bottom half shows the pressure field, with 10 iso-contour equally spaced in $p/\rho U^{2}\in\pm 2$.
@@
 5 shows the time trace of the pressure force acting on the accelerated disk, as well as the equivalent 2D simulation on an accelerating flat plate for comparison.
-The difference in the pressure field is found again in the difference in the forces.
+The impact of the error in in the pressure field is apparent in the difference in the forces.
 Here, the Biot-Savart boundary condition can exactly recover the added-mass force on the first time step, while the reflective boundary condition overestimates it.
 For larger $t^{*}$, the effect of the boundary condition is still observable; the reflective boundary condition significantly overestimates the pressure forces for all $t$.
-Figure 7: Instantaneous pressure force acting on the disk for the two different boundary conditions. The convective time corresponds to the snapshots shown in Fig. 6
+Figure 5: Instantaneous pressure forces normalized by added mass acting on the 2 and 3D disk accelerated from rest for (a) the two different boundary conditions and (b) for different Biot-Savart kernels (tree-sum and $\ell$M) and different subdomain half-width $\tilde{S}$ for the 3D disk.
+The convective time corresponds to the snapshots shown in Fig.
+4.
+Next we demonstrate the low sensitivity of the Eulerian flow solver using the simple partitioned approach of Eq.
+16 to the implementation details of the Biot-Savart kernel summation.
+Fig.
+5(b) compares the instantaneous drag force acting on the 3D disk using the tree-sum and $\ell$M approach over a range of clustering subdomain sizes.
+The results are extremely consistent, with any $\tilde{S}\geq 2$ predicting identical forces and only the smallest possible half-width $\tilde{S}=1$ leading to a few percent increase in the predicted peak force at $t^{*}=2$.
+In addition, the forces are identical using the two approaches despite the additional unpooling step of the $\ell$M.
+This minimal sensitivity means that we can maintain accuracy while maximizing the reconstruction speed.
+For the remainder of the manuscript we will use the $\ell$M approach with $\tilde{S}=2$ for 3D flows, and $\tilde{S}=4$ for 2D.
+We therefore use $O(N)$ operations for clustering the sources and $O(N^{\cal S})$ operations for reconstructing the target velocities for each V-cycle of the pressure solver, which itself is $O(N)$.
+In practice, we find that using Biot-Savart boundary conditions on small domains in 2D can double the simulation speed compared to using reflection condition on the same domain because the artificially fast flow speeds shown in Fig.
+4 decrease the possible time-step.
+In 3D we find simulation times are basically identical using Biot-Savart boundary conditions or reflection conditions on the same domain, with the small reconstruction cost and increased time-steps roughly canceling.
 […]
@@
 We qualitatively compare the shape of the vortex ring forming behind our accelerating square plate with the flow visualization of the same cases in [32], who used dye coloring to visualize the wake forming behind the accelerating plate.
-Results are presented in Figure 9-8 for four instants during and after the acceleration phase.
+Results are presented in Figure 6, 7 for four instants during and after the acceleration phase.
 $a)\,\,t^{*}=2.0$
@@
 […]
-5 Vortex-street applications
-Next, we study the accuracy of the new Biot-Savart conditions on flow cases with vortex wakes that are truncated by the domain boundary, introducing an additional modelling error.
-We use two classical examples of vortex-street flows: (1) the flow around a 2D cylinder at a Reynolds number $\text{Re}=200$, which results in the classical van Kàrmàn wake, and (2) the reversed van Kàrmàn (or propulsive) wake generated by a heaving airfoil in a uniform flow at a range of Strouhal numbers.
-5.1 Flow around a 2D cylinder at $\text{Re}=200$
-We start our analysis with the flow behind a 2D circular cylinder of diameter $D$.
-The cylinder is immersed in a viscous fluid with free-stream velocity $U=(1,0)$.
-The viscosity is set such that the Reynolds number is $\text{Re}=\frac{UD}{\nu}=200$.
-We perform a domain study by varying the domain sizes from $5D\times 3D$ up to $30D\times 24D$ and, for each domain, compare the reflective and Biot-Savart boundary conditions.
-Figure 10 shows the total drag force acting on the cylinder for the various domain sizes and boundary conditions.
-The Biot-Savart boundary conditions yield the best results with an excellent collapse of the drag forcesregardless of the domain size, including the tiny $5Dx3D$ domain which captures no more than two wake vortices at any time.
-Reflective boundary conditions induce extremely large blockage effects for small domains and only slowly converge to the Biot-Savart boundary solution with a domain 48x smaller.
-Additionally, we compare the results obtained with a grid aligned with the mean flow direction to results obtained on a $7D\times 7D$ domain where the flow direction is $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$ with the Biot-Savart boundary conditions.
-Results are also presented in Fig.
-10 and are almost indistinguishable from the results using the aligned domain.
-In addition to the total drag forces acting on the cylinder, we look at the vortex shedding frequency in the wake.
-Here again, the Biot-Savart boundary conditions show an excellent collapse of the shedding frequency, which matches that of classical bluff body vortex shedding frequency of $St\sim 0.3$.
-The high blockage effect induced by the reflective boundary conditions induces fast flow around the cylinder, adding a non-physical domain-size dependance on the shedding frequency.
-Figure 10: (left) Drag force acting on the circular cylinder at $\text{Re}=200$ for various combinations of boundary conditions and domain size.
-The rotated domain consists of a $4D\times 4D$ domain, with the cylinder centered at $1.5D\times 1.5D$ and a free stream velocity of $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$.
-(right) The frequency spectrum of the wake of the cylinder for the same combination of boundary conditions and domain size shows the classical shedding frequency at $St=0.3$.
+4 Developed vortex-wake applications
+Next, we study the accuracy of the new Biot-Savart conditions on flow cases with vortex wakes that are truncated by the domain boundary, introducing an additional modeling error.
+We use two canonical developed vortex-wake examples: (1) the flow around a sphere at $\text{Re}=3700$, and (2) the reversed van Kàrmàn (or propulsive) wake generated by a heaving 2D airfoil over a range of Strouhal numbers.
+4.1 3D flow around a sphere
+Figure 8: Sphere flow vortex wake and drag coefficient at $Re=UR/\nu=3700$ with a resolution of 44 cells per radius.
+(top) Wake visualized by a volumetric rendering of $|\omega|R/U$ at time $tU/R=53$ with the full $3.6D\times 1.5D\times 1.5D$ domain shown.
+Instantaneous (right) and time-averaged (left) drag coefficients for three domains, where $A$ is the domain frontal area.
+The time-averages drag is obtained by averaging the last 100 convective times.
+The flow around a sphere is used to validate the coupled momentum-Biot-Savart update for developed 3D flows.
+A sphere of radius $R$ is immersed in a viscous incompressible flow.
+The Reynolds number is set to $Re=UR/\nu=3700$, we use a resolution of 44 cells per radius and use a domain size of $5C/2\times C\times C$ cells varied over $C\in[128,192,256]$.
+The flow is impulsively started from rest to a uniform velocity $U$.
+We perform the simulations until convective time $tU/R=200$ and Figure 8 shows the drag coefficient when using the Biot-Savart boundary condition is insensitive to the domain size - even for domains only 50% wider than the sphere diameter.
+The averaged drag over the last 100 convective time for all blockage ratios compares favorably to the DNS and LES simulation results obtained in [33, 34].
 4.2 Wake behind a heaving foil
@@
 We next continue our analysis with thrust wakes, where the influence of wake history is expected to be stronger.
-We immerse a rigid foil of length $L$ moving with a linear velocity $\vec{U}=(U,0)$ in a viscous fluid.
-In addition to its steady forward speed, the foil undergoes a pure heave motion of amplitude $h_{0}$
+We immerse a rigid foil of length $L$ moving with velocity $U$ in a viscous fluid. In addition to its steady forward speed, the foil undergoes a pure heave motion of amplitude $h_{0}$
 $y(t)=h_{0}\sin(2\pi ft)$ (24)
@@
 We simulate the flow around this heaving airfoil for 100 convective times to ensure that the wake stabilizes, and we gather pressure and viscous forces for the last 20 convective times.
-Mean lift forces are obtained by time-averaging these forces over the past 20 convective times.
+Mean lift forces are obtained by time-averaging these forces over the last 20 convective times.
 (a) Reflection ($30L\times 20L$)
@@
 (c) Biot-Savart ($6L\times 4L$)
-(d) Biot Savart ($12L\times 8L$)
+(d) Biot-Savart ($12L\times 8L$)
 Figure 9: Snapshot of the deflected wake behind an airfoil at $\text{St}=0.6$ and $\text{Re}=100$ for an amplitude to chord ratio $h_{0}/L=1.0$.
@@
 Vorticity is shown as 10 equally spaced isocontours in the interval $\omega L/U\pm 10$.
-Figure 12: (left) Mean lift force acting on the airfoil at $\text{Re}=100$, for a range of Strouhal numbers for the Biot-Savart and the reflective boundary condition and (right) Poincaré map of the lift and drag coefficients for $St=0.6$
-[…]
-We expect that in any case where wake history effects are highly dominant, the new Biot-Savart boundary condition will require a domain large enough for this wake to develop “stable” dynamics before exiting the domain.
+Figure 10: (left) Mean lift force acting on the airfoil at $\text{Re}=100$, for a range of Strouhal numbers for the Biot-Savart and the reflective boundary condition and (right) Poincaré map of the lift and drag coefficients for $\text{St}=0.6$
+We observe almost no difference between the domain sizes for $\text{St}<0.5$.
+The wake is symmetric, and the airfoil produces zero mean lift forces; Fig.
+10(a).
+For $\text{St}\geq 0.5$, we observe wake deflection due to the high frequency and large circulation vortices pairing up into vortex-couples as seen in Fig.
+9 for $\text{St}=0.6$.
+Fig.
+10(b) shows a Poincaré map for the lift and drag coefficient of the heaving airfoil at $\text{St}=0.6$, with the asymmetry in the lift resulting in a net mean lift coefficient around $C_{L}=-1$ for the very large domain with reflection BCs.
+The new Biot-Savart boundary conditions predict wake deflection for $\text{St}\geq 0.5$ correctly, but truncating this sensitive vortex wake causes the space between the vortex couples to reduce, Fig.
+9(c,d).
+Using a domain extending only four chord lengths behind the foil results in a 10-40% over-prediction of the measured forces, while doubling the domain recovers the correct spacing and forces to within 5%.
+We expect that in any case where wake history effects are highly dominant, the new Biot-Savart boundary condition will require a domain large enough for this wake to develop stable dynamics before exiting the domain.
 However, we should note that the case of 2D deflected thrust wakes is somewhat pathological, depending on long-range vortex interactions only possible in 2D flows.
@@
 5 Conclusion
-In this manuscript, we presented a novel boundary condition formulation for external incompressible flow simulations.
+In this paper, we present a novel coupling to impose unbounded boundary conditions on viscous incompressible flow simulations.
 The method relies on a Biot-Savart integral of the vorticity inside the domain to set the velocity boundary conditions on the domain boundaries.
-We couple these boundary conditions with an explicit projection scheme to solve the momentum and continuity equation in their primitive variables form.
-Leveraging the fast decaying properties of the Biot-Savart kernel, we use a multilevel algorithm to reduce the number of operations required to update the boundary nodes from $O(N)$ to $O(\log N)$ and show that the error introduced by this multilevel approach is bounded and can be estimated.
-We demonstrate that when a body is immersed in a domain using these boundary conditions, a coupling between the immersed body and the domain’s boundary is introduced.
-This coupling only results in a weakly non-linear pressure Poisson equation and a simple differed correction approach can be used to solve the coupled projection/Biot-Savart step, allowing the use of standard Poisson solver.
+We couple these boundary conditions with an explicit projection scheme to solve the momentum and continuity equations in their primitive variable form.
+Critically, we demonstrate that when we apply the Biot-Savart equation to the domain boundaries, a coupling between every point on immersed solid boundaries and the domain boundary is introduced.
+This coupling results in a dense pressure equation that we solve using iterationss on the projection and Biot-Savart update steps, allowing a standard Poisson solver to be used.
+This approach also makes the solution extremely robust, enabling simple and high-speed FMM variants to be used without sacrificing solution accuracy.
+In this work, we use a multi-level clustering method for the $N$ Eulerian-grid vortex sources, and use this to develop a simple Fast Multi-level Method ($\ell$M) for the velocity reconstruction with $O(N)$ complexity that can be run in parallel on CPU and GPU architectures.
+We show that the error introduced by this multilevel approach is bounded and that its influence on the resulting flow field is minimal because of the iterative partitioned approach used for the pressure.
 With different examples of flow with confined vorticity, we show that the novel boundary conditions allow for a significant reduction in the size of the computational domain without any loss in accuracy.
-The method is able to exactly capture added-mass force on an accelerated disk with a computational domain only extending $1/2$ diameter away from the plate.
-Surprisingly, for vortex street applications, the method demonstrates its robustness and perfectly captures the drag forces and vortex shedding frequencies generated by a 2D circular cylinder at $\text{Re}=200$, even with minimal domain sizes ($5D\times 3D$).
-Classical reflection boundary conditions require a domain 48 times bigger to generate time-varying forces indistinguishable from that of the domain with our Biot-Savart boundary conditions.
-Lastly, we show that the method is limited when strong wake history influences the flow in the near wake.
-For a high-amplitude, high-Strouhal number heaving airfoil, we show that capturing the deflected wake of the flow requires significantly larger domain, implying that the computational advantage of the new boundary condition is reduced.
-The generality and accuracy of the new approach, even when simulating flows with vortex wakes that are truncated by the domain boundaries, opens a huge range of potential applications.
-In addition to the welcomed speed-up enabled by using frankly tiny computational domains, the continued accuracy of the simulations challenges the notion of the importance of far-field wake dynamics on the near-body flow, providing compelling evidence to the contrary.
-Appendix A Accounting for symmetries in the Biot-Savart integral
-We briefly describe how symmetries can be accounted for in the multilevel evaluation of the Biot-Savart integral Eq.
+The method exactly captures added-mass force on an accelerated disk with a computational domain only extending $1/2$ diameter away from the plate.
+This domain is also large enough to perfectly replicate the theoretical and previous numerical studies on an impulsively started circular cylinder.
+In comparison, a domain more than 100 times larger is required when using reflection boundary conditions.
+Surprisingly, the method demonstrates its robustness for flows with developed vortex wakes.
+The developed flow on a sphere is found to be largely independent of the domain size, even for blockage ratios greater than 40%.
+Lastly, we show that the method is limited when a strong wake history influences the flow in the near wake.
+For a high-amplitude, high-Strouhal number ($St\geq 0.5$) heaving airfoil, we show that capturing the deflected wake of the flow requires a significantly larger domain, implying that the computational advantage of the new boundary condition is reduced.
+The generality and accuracy of the new approach, even when simulating flows with vortex wakes truncated by the domain boundaries, opens a large range of potential applications.
+In addition to the welcomed speed-up enabled by using minimal computational domains, the continued accuracy of the simulations challenges the notion of the importance of far-field wake dynamics on the near-body flow, providing compelling evidence to the contrary.
+Appendix A Multilevel error analysis
+A.1 Analytic error bound
+We derive the error induced by this multilevel approach, starting with the 2D-case and extending the result to 3D.
+The velocity at location $\vec{x}$ induced by the circulation from a single cell on level $i+1$ is given by (see Fig.
+2)
+$\vec{u}^{(i+1)}(\vec{x})=\frac{\vec{r}}{2\pi|\vec{r}|^{2}}\times\vec{e}_{z}\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}$ (25)
+where $\vec{r}=\vec{y}-\vec{x}$, $y$ is the center of the cell, and $\vec{e}_{z}$ is the out-of-plane unit vector. The velocity induced by the four corresponding cells on level $i$ is
+$\vec{u}^{(i)}(\vec{x})=\mathcal{P}^{(i)\to(i+1)}\left(\frac{\vec{r}_{j}}{2\pi|\vec{r}_{j}|^{2}}\times\vec{e}_{z}\Gamma^{(i)}_{j}\right),$ (26)
+where $\vec{r}_{j}=\vec{y}_{j}-\vec{x}$ and $y_{j}$ are the four cell centers.
+We can further decompose this expression by noting that $\vec{r}_{j}=\vec{y}-\vec{x}-\vec{\delta r}_{j}=\vec{r}-\vec{\delta r}_{j}$.
+For sufficiently large $\vec{r}$ we can assume that $|\vec{r}|\approx|\vec{r}-\vec{\delta r}_{j}|$ which allows us to separate the kernel’s influence into two parts
+$\vec{u}^{(i)}(\vec{x})=\mathcal{P}^{(i)\to(i+1)}\left(\frac{\vec{r}}{2\pi|\vec{r}|^{2}}\times\vec{e}_{z}\Gamma^{(i)}_{j}-\frac{\vec{\delta r}_{j}}{2\pi|\vec{r}|^{2}}\times\vec{e}_{z}\Gamma^{(i)}_{j}\right).$ (27)
+After rearrangement, the first term equals Eq. 25. The second term is the error due to the multilevel approximation. Its magnitude is
+$\varepsilon(\vec{x})^{(i)}=\mathcal{P}^{(i)\to(i+1)}\frac{|\vec{\delta r}_{j}|}{2\pi|\vec{r}|^{2}}\Gamma^{(i)}_{j},$ (28)
+as $\vec{e}_{z}$ is a unit vector.
+On uniform Cartesian meshes, $|\vec{\delta r}_{j}^{(i)}|\sim 2^{l}$, while $|\vec{r}|^{2}\sim(2^{l}\tilde{S})^{2}$.
+An upper bound for the total error in the induced velocity at a point $\vec{x}$ due to the multilevel evaluation of the Biot-Savart integral is therefore
+$\varepsilon(\vec{x})\approx\frac{1}{2\pi\tilde{S}^{2}}\sum_{i=2}^{l}\frac{1}{2^{i}}\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}.$ (29)
+which decays as $\tilde{S}^{-2}$.
+The content of each sum is a decreasing fraction of the total vorticity contained in the domain $\mathcal{D}^{(i)}$.
+As expected, taking $\tilde{S}>\text{Domain}$ yields $\varepsilon(\vec{x})=0$, and as $\Gamma^{(i)}\to 0,\varepsilon(\vec{x})\to 0$, finally as $\tilde{S}\to 0,\varepsilon(\vec{x})\to\infty$.
+Similarly, in 3D the upper bound for the error is given by
+$\varepsilon(\vec{x})\approx\frac{1}{4\pi\tilde{S}^{3}}\sum_{i=2}^{l}\frac{1}{2^{2i}}\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}.$ (30)
+which decays as $\tilde{S}^{-3}$.
+Similarly, the content of the sum is a (even smaller) decreasing fraction of total vorticity contained in each level, improving the method’s accuracy in 3D compared to 2D.
+A.2 2D vortex reconstruction example
+We next use an analytic 2D vortex field to demonstrate the error in the reconstruction method. The velocity field generated by the Lamb-Chaplygin dipole is defined by the scalar stream function
+$\psi(r)=\begin{cases}\frac{-2UJ_{1}(\beta r)}{\beta J_{0}(\beta R)},\quad\text{for}\quad r<R,\\ U\left(\frac{R^{2}}{r}-r\right),\quad\text{for}\quad r\geq R,\end{cases}$ (31)
+where $J_{0}$ and $J_{1}$ are the zeroth and first Bessel functions of the first kind, respectively, and $\beta$ is a parameter with the value $\beta=1.2197\pi/R$.
+The 2D velocity field is given by $\vec{u}=-\vec{e}_{z}\times\nabla\psi$, and the core of the vortex is defined by the isoline $\psi(r=R)=0$.
+We reconstruct the velocity field outside the vortex core using the internal vorticity and our multilevel Biot-Savart method and compare that to the analytic result.
+We reconstruct the velocity on a domain of size $[2^{8},2^{8}]$.
+The dependency of the absolute error in the reconstructed velocity field on the half-width $\tilde{S}$ of the subdomains are shown in Fig.
+11(a).
+Setting $\tilde{S}$ to the domain size $\cal D$ eliminates the source clustering error, but the result is not exact since the analytic vorticity distribution is sampled on a finite grid of size $\Delta x=0.05R$.
+This error depends primarily on the distance from the vortex, Fig.
+11(a).
+Progressively reducing the subdomain size introduces the multilevel error quantified in appendix A, which is only weakly dependent on the distance to the vortex.
+Fig.
+11(b) reports the measured computational speed-up achieved by the multilevel method relative to setting $\tilde{S}=\Omega$, demonstrating up to 300$\times$ acceleration.
+Figure 11: Maximum error in the reconstructed velocity field generated by a 2D Lamb-Chaplygin vortex dipole for various kernel sizes and distances from the vortex core (left).
+Corresponding speed-up obtained by reducing the kernel size for the multilevel algorithm (right).
+The mesh has dimensions $[2^{8},2^{8}]$.
+A.3 3D vortex reconstruction example
+This validation is repeated in 3D using the classical Hill vortex, whose radial and tangential velocity components are given by
+$v_{r}$ $=\frac{1}{r^{2}\sin\theta}\frac{\partial\psi}{\partial\theta},\qquad v_{\theta}=-\frac{1}{r\sin\theta}\frac{\partial\psi}{\partial r}.$ (32)
+and the stream function
+$\psi(r,\theta)=\begin{cases}-\frac{3U}{4}\left(1-\frac{r^{2}}{R^{2}}\right)r^{2}sin^{2}\theta\quad\text{for}\,r\leq R,\\ \,\,\frac{U}{2}\left(1-\frac{r^{3}}{R^{3}}\right)r^{2}sin^{2}\theta\quad\text{for}\,r>R.\end{cases}$ (33)
+with the corresponding errors and speed-up shown in Figure 12.
+Note that the analytic vorticity of the hill vortex is discontinuous, increasing the truncation error near the vortex compared to the smooth 2D Lamb dipole.
+The multilevel error decreases extremely quickly as $\tilde{S}$ increases for three-dimensional flows.
+This is due to the scaling of the Biot-Savart kernel itself.
+Similarly, the speed-up is much faster in 3D because each doubling gathers 8 cells instead of 4, with a maximum observed speed-up of nearly 20,000$\times$ using the multilevel method.
+Figure 12: Maximum error in the reconstructed velocity field generated by a 3D Hill vortex for various kernel sizes and distances from the vortex core (left).
+Corresponding speed-up obtained by reducing the kernel size for the multilevel algorithm (right).
+The mesh has dimensions $[2^{8},2^{8},2^{8}]$.
+Appendix B Accounting for symmetries in the Biot-Savart integral
+Symmetries can be accounted for in the multilevel evaluation of the Biot-Savart integral Eq.
 8.